### v1

# MPCache: MPC-Friendly KV Cache Eviction for Efficient Private Large Language Model Inference

Abstract
Private large language model (LLM) inference based on secure multi-party computation (MPC) offers cryptographically-secure protection for both user prompt and proprietary model weights. However, it suffers from large latency overhead especially for long input sequences. While key-value (KV) cache eviction algorithms have been proposed to reduce the computation and memory cost for plaintext inference, they are not designed for MPC and cannot benefit private inference easily. In this paper, we propose an accurate and MPC-friendly KV cache eviction framework, dubbed MPCache. MPCache is built on the observation that historical tokens in a long sequence may have different effects on the downstream decoding. Hence, MPCache combines a look-once static eviction algorithm to discard unimportant tokens and a query-aware dynamic selection algorithm to further select a small subset of tokens for attention computation. As existing dynamic selection algorithms incur too much latency, we propose a series of optimizations to drastically reduce the KV cache selection overhead, including MPC-friendly similarity approximation, hierarchical KV cache clustering, and cross-layer index sharing strategy. With extensive experiments, we demonstrate that MPCache consistently outperforms prior-art KV cache eviction baselines across different LLM generation tasks and achieves 1.82.01×1.8$\sim$ 2.01$\times$1.8 ∼ 2.01 × and 3.398.37×3.39$\sim$ 8.37$\times$3.39 ∼ 8.37 × decoding latency and communication reduction on different sequence lengths, respectively.
Large Language Models, KV Cache Compression, Private Inference, Multi-party Computation
1 Introduction
Large language models (LLMs) have recently demonstrated remarkable ability in a wide range of applications such as document summarization (Huang et al., 2021; Narayan et al., 2018; Zhang et al., 2024b), question answering (Kočiskỳ et al., 2018; Dasigi et al., 2021; Yang et al., 2018), and dialogue systems (Thoppilan et al., 2022; Chiang et al., 2023; Taori et al., 2023). However, LLM-based machine learning as a service (MLaaS) on the cloud has raised serious privacy concerns as the users are required to upload their prompts to the cloud, which may contain sensitive personal information. Meanwhile, the service provider is unwilling to offload the trained model to the user to protect the proprietary model weights. Secure multi-party computation (MPC)-based private inference has been proposed to address the privacy concerns (Goldreich, 1998; Mohassel & Rindal, 2018; Huang et al., 2022; Rathee et al., 2020; Gupta et al., 2023). MPC enables the users and the cloud to conduct the LLM inference jointly, but nothing else can be derived beyond the final inference results.
However, MPC-based LLM inference faces serious efficiency challenges, especially for long input sequences. We profile the decoding efficiency of GPT-2 with the Secretflow framework (Ma et al., 2023) using recent 2-party computation (2PC) (Lu et al., 2023) and 3-party computation (3PC) protocols (Dong et al., 2023). As can be observed in Figure 1(a) and (b), attention dominates the latency and communication for both 2PC and 3PC protocols. Moreover, Softmax accounts for the majority of the overall cost due to its complex operators including max, exponential, and division, especially with an increasing sequence length.
Figure 1: (a) Breakdown of decoding latency and communication for one token generation with a sequence length of 512. Attention dominates the latency and communication for both 3PC and 2PC protocols. (b) The cost of Softmax scales with the sequence length rapidly. (c) Inference cost before and after KV cache eviction. Blocks in slash indicate the extra overhead introduced by eviction.
To reduce the cost of private LLM inference, previous works focus on developing more efficient MPC protocols (Lu et al., 2023; Dong et al., 2023; Pang et al., 2023; Hou et al., 2023), replacing non-linear activation functions with more MPC-friendly operators (Liu & Liu, 2023; Li et al., 2022; Zeng et al., 2023), or directly modifying the model architecture (Rathee et al., 2024). However, they still incur significant overhead or require expensive finetuning or re-training, and cannot be directly applied to LLMs. Another line of works leverages key-value (KV) cache eviction to reduce the number of tokens involved in the attention computation (Zhang et al., 2024f; Ge et al., 2023; Liu et al., 2024c; Zhao et al., 2024b; Zhang et al., 2024d; Fu et al., 2024). Although they have demonstrated significant memory and computation reduction for plaintext LLM inference without the need of finetuning, they are not MPC-friendly. As shown in Figure 1(c), directly applying an existing KV cache eviction algorithm (Liu et al., 2024b) incurs even more communication and latency overhead over the baseline model since it introduces expensive operators in MPC, including top-$k$ ranking, token gathering, etc, as elaborated in Section 3. Therefore, there is an urgent need for a training-free MPC-friendly KV cache eviction algorithm to improve the efficiency of private LLM inference.
To overcome the heavy overhead of attention computation, we make the following observations that motivate our MPCache: 1) LLM attention maps are overall sparse for long input prompts, motivating us to perform static eviction and directly prune the KV cache of unimportant tokens; 2) LLM attention shows dynamism and token-wise locality (Liu et al., 2023), motivating us to build an efficient clustering algorithm for dynamic selection of the KV cache; 3) LLM attention of adjacent layers shows similar patterns, motivating us to share the KV cache selection for adjacent layers to further improve efficiency. Our contributions can be summarized as follows:
  • We observe the cost of MPC-based LLM inference mainly comes from attention computation and propose MPCache, an MPC-friendly KV cache eviction framework to reduce the LLM inference latency and communication.
  • We identify the challenges when applying KV cache eviction to MPC. To tackle the problems, MPCache combines look-once static eviction and query-aware dynamic selection with a series of optimizations, including MPC-friendly similarity approximation, hierarchical KV cache clustering, and cross-layer index sharing strategy.
  • With extensive experiments, we demonstrate the performance of MPCache consistently exceeds the prior-art KV cache eviction algorithms across different generation tasks and achieves up to 2.01×2.01$\times$2.01 × and 8.37×8.37$\times$8.37 × decoding latency and communication, respectively.
2 Problem Formulation and Background
2.1 Problem Formulation
Algorithm 1 Formulation of KV cache eviction for one layer
Input :
Query, key, and value cache $\mathbf{q}\in\mathbb{R}^{H\times 1\times d},\mathbf{K,V}\in\mathbb{R}^{H\times T\times d}$, where $T,H,d$ denote the sequence length, number of heads, and embedding dimension.
Output :
Sparse attention output $\mathbf{O}\in\mathbb{R}^{H\times 1\times d}$.
$\mathbf{sim}=\mathrm{Sim}(\mathbf{q},\mathbf{K})$; $\triangleright$ Similarity approximation
$\textbf{indices}=\mathrm{topk}(\mathbf{sim},k=k)$; $\triangleright$ Top-$k$ ranking
$\mathbf{K}^{\prime}=\mathbf{K}.\mathrm{gather[\textbf{indices}]},~{}\mathbf{V}^{\prime}=\mathbf{V}.\mathrm{gather[\textbf{indices}]}$; $\triangleright$ Token gathering based on indices
$\mathbf{O}=\mathrm{Softmax}(\mathbf{q}\cdot\mathbf{K}^{\prime\top}/\sqrt{d})\cdot\mathbf{V}^{\prime}$; $\triangleright$ Sparse attention
return $\mathbf{O}$.
Generative LLM inference can be divided into prefill and decoding stages (refer to Appendix A). We formally describe the decoding process with KV cache eviction in Algorithm 1. The KV cache eviction policy, denoted as $\mathcal{P}$, aims to minimize the attention computation by only preserving a subset of tokens, which typically involves three steps: 1) $\mathcal{P}$ first computes the similarity between the query and key cache of previous tokens (line # 1); 2) $\mathcal{P}$ then ranks the previous tokens based on the similarity score and applies the top-$k$ algorithm to determine the indices of relevant tokens (line # 2); 3) the KV cache is then retrieved based on the indices, denoted as token gathering (line # 3) [^1] We describe the procedure and protocol of token gathering in Appendix B.4., followed by sparse attention computation with the selected KV cache (line # 4). To compute the similarity in line # 1, existing works have used accumulated attention score of the historical tokens (Liu et al., 2024c; Zhang et al., 2024f; Zhao et al., 2024b; Yang et al., 2024; Zhang et al., 2024d) or cosine similarity (Liu et al., 2024b; Xiao et al., 2024a). KV cache eviction reduces the attention computation complexity from $\mathcal{O}(Td)$ to $\mathcal{O}(kd)$, where $T,d$ denote the sequence length and embedding dimension, respectively, and $k\ll T$. However, it introduces MPC-unfriendly operations, including similarity approximation, top-$k$ ranking, and token gathering, hindering its benefits in MPC-based LLM inference. Hence, the goal of this work can be summarized as follows:
“How can we design an MPC-friendly KV cache eviction algorithm $\mathcal{P}$ to minimize MPC-based LLM inference overhead without sacrificing LLM performance?”
2.2 Background
Related works. There has been a surge in improving the efficiency of private LLM inference. Existing works focus on the protocol optimization (Pang et al., 2023; Dong et al., 2023; Lu et al., 2023; Hou et al., 2023) or replacing non-linear functions with MPC-friendly operators (Liu & Liu, 2023; Li et al., 2022; Zeng et al., 2023; Mishra et al., 2020; Dhyani et al., 2023). However, they either incur large overhead for long sequences or require expensive training. KV cache eviction has been widely explored for plaintext inference and can be mainly classified into 3 categories: 1) fixed-pattern algorithms like Xiao et al. (2023); Beltagy et al. (2020); Zaheer et al. (2020) always keep the tokens in the same position, e.g., initial and recent tokens across decoding steps, lacking flexibility for different LLMs and contexts; 2) static algorithms like Zhang et al. (2024f); Zhao et al. (2024b); Zhang et al. (2024d); Li et al. (2024b); Ge et al. (2023) discard tokens based on the accumulated attention scores of historical tokens, which are efficient as the eviction is usually only conducted once but suffers from large performance degradation when the compression ratio is high; 3) dynamic algorithms like Xiao et al. (2024a); Tang et al. (2024b); Liu et al. (2024b); Hooper et al. (2024a) compute the similarity between the current query and key cache for every decoding step, which is more accurate but requires repetitive selection at each step. Recently, there are also some learning-based methods (Xiao et al., 2024b; Gao et al., 2024) which requires extra training. Different from prior works in Table 1, MPCache is a training-free framework that combines static and dynamic algorithms, and leverages hierarchical clustering with a series of MPC-friendly optimizations, simultaneously achieving high efficiency and performance. We leave a detailed review of existing works in Appendix A.
Table 1: Qualitative comparison with prior works of attention optimization.
Representative
Work
Method
Similarity
Approximation
Top-$k$
Ranking
Token
Gathering
Cross-layer
Optimization
MPC
Efficiency
Model
Performance
Li et al. (2022)
Non-linear
Approximation
Finetuning
Required
Not Applied
to LLM
Xiao et al. (2024b) Learning-based - - - -
Training
Required
High
Xiao et al. (2023) Fixed-pattern - - - - High Low
Li et al. (2024b) Static Accumulated Attn. Score Usually Once Token-wise - High Low
Liu et al. (2024b) Dynamic Token-wise Similarity Token-wise per Step Token-wise - Low High
MPCache (ours) Static+Dynamic
Hierarchical Clustering,
Cluster-wise Similarity
Paralleled Computation,
Cluster-wise per Step
Cluster-wise
Adjacent Layer
Index Sharing
High High
MPC preliminaries. MPC (Goldreich, 1998) is a cryptographic technique recently developed and leveraged to enable LLM inference while protecting the privacy of both data and model. In an MPC framework, to protect a certain tensor, it is often split into multiple secret shares and distributed across different parties involved in the computation (Lu et al., 2023; Dong et al., 2023; Mohassel & Rindal, 2018). Dedicated protocols have been developed to support LLMs’ linear and non-linear operations (Lu et al., 2023; Pang et al., 2023; Dong et al., 2023).
Threat model. In this work, we adopt an honest-but-curious threat model and apply MPCache to both 2PC and 3PC protocols, which involve 2 parties and 3 parties in the computation, respectively. Following Li et al. (2022); Zeng et al. (2023); Li et al. (2024a), MPCache is built upon existing cryptographic primitives and focuses on optimizing the LLM inference algorithm. The security can hence be guaranteed. Our threat model assume all the parties are aware of the model architecture and number of pruned tokens, which is consistent with Zhang et al. (2024e); Li et al. (2024a); Kundu et al. (2023); Cho et al. (2022). We argue that this information does not compromise the client’s data or inference results, nor does it enable the client to access the model’s parameters. We refer interested readers to Appendix B, where the threat model and underlying protocols are more clearly explained.
3 Motivations and Challenges
In this section, we discuss the key observations that motivate the design of MPCache.
Figure 2: (Upper) token types in attention maps where means the token is selected and means the token is not selected. (Lower) three types can be observed in the attention map with more tokens.
Observation 1: the attention map of a long input sequence is usually sparse, and the KV cache of historical tokens demonstrates different impacts over the downstream decoding. We show the attention map of different heads and layers of LLaMA-2-7B in Figure 2 and leave visualizations of larger attention maps in Appendix C. From Figure 2, we can classify different tokens into 3 categories: 1) important to all tokens (IA in red box): the attention scores remain high for the entire column, e.g., 0th and 1st columns in Figure 2(a), indicating these tokens are important for the generation of all downstream tokens and hence, need to be always preserved; 2) un-important to all tokens (UIA in blue box): the attention scores remain low for the entire column, e.g., 2nd and 3rd columns in Figure 2(a), indicating these tokens can be discarded without impacting the downstream decoding; 3) important to certain tokens (IC in orange box): the attention scores vary for different tokens, e.g., 4th and 5th columns in Figure 2(a), indicating these tokens impact a subset of downstream tokens, and hence, cannot be directly pruned.
We verify the observation on LLaMA-2-7B with different input sequence lengths. As shown in Figure 3(a), almost 60% tokens can be statically evicted while preserving the LLM performance. While further pruning the remaining tokens starts to degrade the LLM performance, as shown in Figure 3(b), in each decoding step, only less than 20% of the remaining tokens contribute to the decoding. The above observation motivates us to statically evict the KV cache of UIA tokens and dynamically select a subset of IC tokens at each decoding step.
Figure 3: Motivating inspirations of MPCache. (a) Statically evicting 60% tokens during the prefill stage still maintains the performance; (b) less than 20% tokens contribute to token decoding; (c) cross-layer commonality among different numbers of adjacent layers.
Observation 2: dynamic KV cache selection incurs non-negligible overhead in MPC. While dynamic KV cache selection is accurate and reduces the attention computation cost, it incurs non-negligible overhead due to MPC-unfriendly operations. In Figure 1(c), we show the extra overhead when 5% tokens are dynamically selected. The MPC-unfriendly operations mainly include:
  • Similarity computation (Algorithm 1 line # 1): cosine similarity is widely used, which requires computing the multiplication between the current query with the key cache of all previous tokens;
  • Top-$k$ ranking (Algorithm 1 line # 2): to compute the indices of relevant tokens, top-$k$ is usually inevitable (Zhang et al., 2024f; Ge et al., 2023; Zhao et al., 2024b). Unlike plaintext inference, top-$k$ ranking in MPC involves frequent comparison protocol, which incurs high latency and communication cost (Rathee et al., 2020).
  • Token gathering (Algorithm 1 line # 3): after the top-$k$ ranking, the KV cache of selected tokens is gathered based on the indices. Unlike plaintext inference, such gathering protocol in MPC is much more expensive since both KV cache and indices are ciphertexts. Therefore, as described in Algorithm 2, each index is first converted to a one-hot vector and then multiplied with the KV cache, requiring repetitive MPC-unfriendly comparison protocols.
Inspired by token-wise locality (Liu et al., 2023; Zhu et al., 2023), our key insight is to group the adjacent tokens into clusters, which can reduce the complexity of dynamic selection in proportion to the cluster size. However, this introduces extra questions on how to measure the similarity between a cluster and the current query, how to build the cluster, etc, which is discussed in Section 4.3.
Observation 3: adjacent layers share similar top-$k$ ranking of KV cache, providing an extra opportunity for efficiency optimization. Due to the residual, we hypothesize adjacent layers may share a similar top-$k$ ranking of the KV cache. To verify the assumption, we define commonality score to measure the ratio of common top-$k$ indices of $m$ adjacent layers as below:
$\frac{1}{k(L-m)}\sum_{l=1}^{L-m}\left|\bigcap_{i=l}^{l+m}\mathbf{idx}_{i}[:k]\right|,$
[…]
4 MPCache: An MPC-friendly LLM Inference Framework
4.1 Overview of MPCache
Figure 4: Overview of our proposed MPCache.
Framework. Driven by the observations, we propose an MPC-friendly KV cache eviction framework, dubbed MPCache. The overview is shown in Figure 4, and it consists of two steps: 1) look-once static eviction during the prefill stage to discard the UIA tokens (Section 4.2); 2) query-aware dynamic selection during the decoding stage to select only a small subset of the remaining IC tokens for sparse attention (Section 4.3). A series of MPC-friendly optimizations are proposed to reduce the overhead of selection (Section 4.3 and 4.4). The pseudocode is shown in Algorithm 3 in Appendix D and the overall data flow is shown in Appendix H.
Symbol definition. For clarity, we summarize the symbols used in this paper. We define $L$ as the number of layers, $H$ as the number of attention heads, $T$ as the number of tokens, $d$ as the embedding dimension, $s$ as the cluster size, and $C$ as the number of clusters.
4.2 Step 1: Look-once Static KV Cache Eviction Algorithm
To prune the KV cache of UIA tokens as observed in Section 3, we use static eviction during the prefill stage. To measure the token importance and identify UIA tokens, we compute the attention map and then, accumulate the attention scores for each token. Similar to Zhang et al. (2024f); Liu et al. (2024c); Li et al. (2024b); Zhao et al. (2024b), we find it is sufficient to only sum up the scores of the last 20% tokens in the prompt. Then, we rank the accumulated attention scores to select the important tokens with the highest scores and discard the rest UIA tokens. The algorithm is only performed once.
Figure 5: The illustration of static eviction.
Protocol complexity analysis. Compared to the baseline computation of the prefill stage, static eviction only involves accumulating the attention scores, which are local without any communication, and a top-$k$ ranking. Because the static eviction is performed only once, the cost can be amortized by the entire generation process, and hence, becomes negligible. Meanwhile, with UIA tokens pruned, the efficiency of the dynamic selection can be improved for all decoding steps. Hence, the static eviction algorithm helps improve the overall efficiency.
4.3 Step 2: MPC-friendly Dynamic KV Cache Selection Algorithm
Figure 6: Hierarchical and dynamic KV cache clustering and selection procedure.
To reduce the overhead of dynamic token selection as shown in Figure 1(c), we propose to group the KV cache of adjacent tokens into clusters as shown in Figure 6(b). The most important question is “how to aggregate the information of a cluster and measure the importance of each cluster accurately and efficiently?”
MPC-friendly similarity approximation with clustering. A naive method for similarity approximation is to compute the average of the key cache within a cluster and directly compute the cosine similarity with the average. However, as shown in Figure 7, the naive approach incurs large performance degradation. Our intuition is that the approximation should preserve the impact of important tokens as much as possible. Hence, we use the maximum dot product between the query and the key cache cluster. Specifically, given a query $\mathbf{q}\in\mathbb{R}^{1\times d}$, a key cache cluster of $s$ tokens $\mathbf{K}_{c}\in\mathbb{R}^{s\times d}$, the similarity can be designed as
$\mathrm{Sim}(\mathbf{q},\mathbf{K}_{c})=\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}\cdot\mathbf{k}=\max_{\mathbf{k}\in\mathbf{K}_{c}}\sum_{i=0}^{d-1}\mathbf{q}_{i}\mathbf{k}_{i}\leq\sum_{i=0}^{d-1}\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}_{i}\mathbf{k}_{i}.$
We obtain the upper bound of similarity, and further have
$\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}_{i}\mathbf{k}_{i}=\begin{cases}\mathbf{q}_{i}\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}&\mathrm{if}~{}\mathbf{q}_{i}\geq 0,\\ \mathbf{q}_{i}\min_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}&\mathrm{if}~{}\mathbf{q}_{i}<0.\end{cases}$
Define $\mathbf{r}^{\max}$ and $\mathbf{r}^{\min}$, where $\mathbf{r}_{i}^{\max}=\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}$ and $\mathbf{r}_{i}^{\min}=\min_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}$. Then, we have
$\mathrm{Sim}(\mathbf{q},\mathbf{K}_{c})\leq\sum_{i=0}^{d-1}\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}_{i}\mathbf{k}_{i}=\sum_{i=0}^{d-1}\max(\mathbf{q}_{i}\mathbf{r}^{\max}_{i},\mathbf{q}_{i}\mathbf{r}^{\min}_{i}).$ (1)
Protocol complexity analysis. During the decoding stage, $\mathbf{r}_{i}^{\max}$ and $\mathbf{r}_{i}^{\min}$ of each cluster only need to be computed once. Hence, the computation cost can be amortized and become negligible. However, for each decoding step, we still need to compute $\mathcal{O}(LCd)$ multiplications, i.e., $\mathbf{q}_{i}\mathbf{r}^{\max}_{i}$ and $\mathbf{q}_{i}\mathbf{r}^{\min}_{i}$, as well as $\mathcal{O}(LCd)$ $\max$ operations in Equation (1), which still incur non-negligible overhead.
Linearization and reordering. To avoid the MPC-unfriendly $\max$ operation in Equation (1), we further propose to approximate the similarity as below:
$\mathrm{Sim}(\mathbf{q},\mathbf{K}_{c})\approx\sum_{i=0}^{d-1}\alpha\cdot\mathbf{q}_{i}\mathbf{r}^{\max}_{i}+(1-\alpha)\cdot\mathbf{q}_{i}\mathbf{r}^{\min}_{i},$
where $\alpha\in[0,1]$ is a hyperparameter. As can be observed, when $\alpha=1$, $\mathbf{q}_{i}\mathbf{r}^{\max}_{i}$ is always selected while $\mathbf{q}_{i}\mathbf{r}^{\min}_{i}$ is always selected when $\alpha=0$. After the linearization, there is an opportunity to further reduce the multiplications by reordering the computation as
$\sum_{i=0}^{d-1}\alpha\cdot\mathbf{q}_{i}\mathbf{r}^{\max}_{i}+(1-\alpha)\cdot\mathbf{q}_{i}\mathbf{r}^{\min}_{i}=\sum_{i=0}^{d-1}\mathbf{q}_{i}\cdot(\alpha\mathbf{r}^{\max}_{i}+(1-\alpha)\mathbf{r}^{\min}_{i}).$ (2)
$\alpha\mathbf{r}_{i}^{\max}$ and $(1-\alpha)\mathbf{r}_{i}^{\min}$ are first added up without introducing extra communication, and the multiplication with $\mathbf{q}_{i}$ is reduced by 2$\times$. Compared with the maximum dot product in Figure 7, our method significantly reduces the cost while maintaining the performance. We empirically choose $\alpha=0.6$, and leave more discussions to Appendix F and a theoretical analysis to Appendix G.
Protocol complexity analysis. MPCache reduces the number of $\max$ operations from $\mathcal{O}(LCd)$ to 00 and reduce the multiplication complexity by 2$\times$. It is worth noting that clustering also benefits the token gathering protocol: 1) the number of comparisons in one-hot vector conversion is reduced by $\frac{T}{C}\times$; 2) the bit width of one-hot vector is reduced by $\frac{\log T}{\log C}\times$. Table 2 shows an example of selecting top-25% tokens with $T=1024,C=64$, and can be observed that the overhead is drastically reduced.
Figure 7: Comparison among maximum dot product (MP), average, and our method with different $\alpha$’s on TriviaQA.
Table 2: The complexity analysis of token gathering protocol where $k_{1}=0.25T,k_{2}=0.25C$.
Bit Width # Comparison Lat. Comm. Example Lat. Example Comm.
Baseline Protocol $\log T$ $T$ $\mathcal{O}(T\log T)$ $\mathcal{O}(k_{1}T\log T)$ 4.780s 416.0MB
MPCache (ours) $\log C$ $C$ $\mathcal{O}(C\log C)$ $\mathcal{O}(k_{2}C\log C)$ 0.065s 1.125MB
Improvement $\frac{\log T}{\log C}\times$ $\frac{T}{C}\times$ $\frac{T\log T}{C\log C}\times$ $\frac{k_{1}T\log T}{k_{2}C\log C}\times$ 73.5$\times$ 369.8$\times$
Hierarchical KV cache clustering. Another question is “how to build the KV cache cluster efficiently?” Since larger cluster sizes have higher efficiency at the cost of worse performance, our key insight is to trade off the selection overhead and model performance. Inspired by hierarchical reinforcement learning (Xu et al., 2023), we propose to cluster the KV cache of adjacent tokens with a hierarchical structure as shown in Figure 6(a) that performs coarse-grained (larger cluster size) to fine-grained (smaller cluster size) selection progressively. Generally, we divide the KV cache into $n$ levels and progressively select the clusters level by level. Then, at the fine-grained level, we only need to select from the remaining clusters, thereby reducing the selection complexity. Hierarchical structure, including the cluster size and selection ratios at different levels, can influence the performance-efficiency trade-off, which is discussed in Section 5.4.
4.4 Cross-layer Index Sharing
Figure 8: Commonality score between two adjacent layers on LLaMA-2-7B.
To leverage the observation that adjacent layers share similar top-$k$ ranking of KV cache, we propose a cross-layer index sharing strategy that enables adjacent layers to share the same selected token indices to further reduce the cost of dynamic selection. Since two adjacent layers show the highest commonality score in Figure 3(c), we choose to share the indices between two adjacent layers. In Figure 8, we observe the first two layers have a low commonality score while other layers have higher scores due to the residual, so we do not apply sharing to the first two layers. Cross-layer index sharing effectively reduces the extra overhead introduced by dynamic selection. We discuss how the number of adjacent layers affects the trade-off in Section 5.4.
5 Empirical Evaluation
Figure 9: Comparison with fixed-pattern and static KV cache eviction algorithms.
5.1 Experimental Setups
Models and datasets. Our experiments are based on LongChat-7B-V1.5-32K (Li et al., 2023) on LongBench (Bai et al., 2023) [^2] https://github.com/THUDM/LongBench: HotpotQA (Yang et al., 2018), NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), GovReport (Huang et al., 2021), TriviaQA (Joshi et al., 2017), and PassageRetrieval (Bai et al., 2023). We also apply our method to LLaMA-2-7B/13B (Touvron et al., 2023) on 5-shot XSUM (Narayan et al., 2018) and LLaMA-3-8B-Instruct (Dubey et al., 2024) on LongBench. To save GPU memory when processing long contexts, we leverage FlashAttention (Dao et al., 2022) during the prefill stage.
Baselines. For comparison, we choose prior-art static and dynamic KV cache eviction baselines, including H2O (Zhang et al., 2024f), StreamingLLM (Xiao et al., 2023), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024b), InfLLM (Xiao et al., 2024a), and LongCache (Liu et al., 2024b). Detailed descriptions of the baselines and our setups can be found in Appendix F.
Experimental environment. For performance evaluation, our experiments are conducted based on LongBench on an NVIDIA A100 80GB GPU. For efficiency evaluation, our experiments are based on Secretflow [^3] https://github.com/secretflow/spu (SPU V0.9.1) (Ma et al., 2023) and follow the protocols of PUMA (Dong et al., 2023). We optimize the top-$k$ protocol in Secretflow with parallelization to support efficient head-wise KV cache selection (see Appendix B for details). The latency is evaluated under the LAN setup (Rathee et al., 2020). We evaluate the efficiency using the architecture of GPT-2 and LLaMA-2, and since securely evaluating a full-size 7B model in SPU exceeds our hardware resources, we set a smaller embedding dimension of 1024 in our evaluation.
5.2 Performance Evaluation
In Figure 9 and Table 4, we comprehensively compare MPCache with prior-art KV cache eviction methods and make the following observations: 1) comparison with fixed-pattern and static algorithms. MPCache consistently outperforms prior-art methods, including H2O, StreamingLLM, TOVA, and SnapKV across different datasets. These methods statically discard the tokens while MPCache dynamically selects a subset of tokens based on the current queries. MPCache shows decent scalability to different KV cache budgets. For example, on HotPotQA and NarrativeQA, MPCache achieves comparable performance as full cache, even only $\sim$5KV cache preserved; 2) comparison with dynamic algorithms. MPCache achieves comparable and even better performance compared with InfLLM and LongCache. On NarrativeQA, MPCache achieves 1.32$\times$ and 2.39$\times$ latency reduction with a higher F1 score compared with InfLLM and LongCache, respectively; 3) scalability of MPCache. We extend the idea of combining static and dynamic algorithm to LLaMA-2-13B in Table 3 and LLaMA-3-8B-Instruct in Table 5, demonstrating the superior performance of MPCache.
Table 3: Comparison of different parameter scales on XSUM.
Budget 10% 5%
[…]
Full Cache 11.90 13.60 11.90 13.60
H2O 10.50 13.24 4.886 9.081
MPCache (ours) 11.10 13.44 10.08 13.08
Table 4: Comparison with dynamic KV cache eviction algorithms on different datasets and budgets. “(a×$\times$)” means MPCache achieves a×$\times$ efficiency improvement compared with the baselines.
Dataset Cache Budget InfLLM LongCache MPCache (ours)
Perf. (%$\uparrow$ Lat. (s)$\downarrow$ Perf. (%$\uparrow$ Lat. (s)$\downarrow$ Perf. (%$\uparrow$ Lat. (s)$\downarrow$
HotpotQA Full 31.16 75.52 31.16 75.52 31.16 75.52
5% 28.20 51.64 (1.30$\times$) 24.31 89.46 (2.24$\times$) 30.27 39.85
10% 29.01 68.04 (1.28$\times$) 24.69 123.1 (2.30$\times$) 30.05 53.32
TriviaQA Full 82.67 75.52 82.67 75.52 82.67 75.52
5% 75.65 51.64 (1.38$\times$) 59.85 89.46 (2.39$\times$) 75.61 37.37
10% 82.75 68.04 (1.34$\times$) 60.56 123.1 (2.43$\times$) 82.45 50.75
NarrativeQA Full 17.02 75.52 17.02 75.52 17.02 75.52
5% 12.80 47.74 (1.32$\times$) 14.65 86.42 (2.39$\times$) 17.23 36.13
10% 13.74 63.49 (1.28$\times$) 15.69 121.4 (2.45$\times$) 17.35 49.46
PassageRetrieval Full 32.50 75.52 32.50 75.52 32.50 75.52
5% 6.161 51.64 (1.15$\times$) 21.42 89.46 (1.99$\times$) 19.75 44.82
10% 8.872 68.04 (1.16$\times$) 24.92 123.1 (2.10$\times$) 27.75 58.47
Qasper Full 27.58 75.52 27.58 75.52 27.58 75.52
8% 20.53 64.52 (1.45$\times$) 24.53 136.9 (3.08$\times$) 23.86 44.39
16% 23.90 72.84 (1.33$\times$) 26.07 225.9 (4.12$\times$) 24.95 54.77
Table 5: Extension to LLaMA-3-8B-Instruct on with an average KV cache size of 2048.
Method
Qasper
(F1 Score)
MultiFieldQA
(F1 Score)
HotpotQA
(F1 Score)
2WikiMultihopQA
(F1 Score)
MuSique
(F1 Score)
TriviaQA
(F1 Score)
TREC
SAMSum
(Rouge-L)
Full Cache 29.75 41.12 45.55 35.87 22.35 90.56 73.0 41.88
SnapKV 25.78 38.13 40.12 32.01 16.86 83.22 70.0 31.75
H2O 26.85 39.54 44.30 32.92 21.09 90.56 53.0 41.84
MPCache (ours) 29.45 40.30 44.32 35.91 22.66 90.43 73.0 42.42
5.3 Inference Efficiency Evaluation
In Figure 10, we benchmark the decoding efficiency with different sequence lengths ranging from 512 to 2048. We compare MPCache with model without KV cache, with full KV cache, LongCache, and head merging (Rathee et al., 2024; Bian et al., 2021). From the results, we make the following observations: 1) KV cache is crucial for private LLM inference since it avoids re-computation of the KV cache of the previous tokens. As shown in the purple bar, the overhead increases by hundreds of times compared with using the KV cache; 2) compared with full KV cache on LLaMA-2, MPCache achieves 1.592.01×1.59$\sim$ 2.01$\times$1.59 ∼ 2.01 ×, 1.461.97×1.46$\sim$ 1.97$\times$1.46 ∼ 1.97 ×, and 1.261.8×1.26$\sim$ 1.8$\times$1.26 ∼ 1.8 × latency reduction and 3.394.18×3.39$\sim$ 4.18$\times$3.39 ∼ 4.18 ×, 4.335.92×4.33$\sim$ 5.92$\times$4.33 ∼ 5.92 ×, and 5.518.37×5.51$\sim$ 8.37$\times$5.51 ∼ 8.37 × communication reduction with different sequence lengths, respectively; 3) compared with LongCache which dynamically selects tokens without static eviction and clustering, MPCache even achieves 3.85$\times$ and 19.47$\times$ latency and communication reduction, respectively. We further discuss 2PC protocol (Lu et al., 2023) in Appendix F.
5.4 Ablation Studies of MPCache
Figure 10: Evaluation on per-token generation latency and communication.
Figure 11: Step-by-step ablation study of MPCache.
Table 6: Comparisons among different hierarchical structures.
Level 1 (Coarse-grained)
Level 2 (Fine-grained)
F1 Score (%) Comm. (MB)
$s32~{}(0.9)$ $s16~{}(0.22)$ 29.6 163.5
$s32~{}(0.7)$ $s16~{}(0.28)$ 30.1 144.0
$s32~{}(0.5)$ $s16~{}(0.40)$ 30.2 140.2
$s32~{}(0.3)$ $s16~{}(0.67)$ 29.2 108.8
$s64~{}(0.9)$ $s16~{}(0.22)$ 29.5 158.1
$s64~{}(0.7)$ $s16~{}(0.28)$ 29.3 110.1
$s64~{}(0.5)$ $s16~{}(0.40)$ 29.1 104.9
$s64~{}(0.3)$ $s16~{}(0.67)$ 29.0 69.12
Effectiveness of different optimizations. In Figure 11, we demonstrate the effectiveness of our proposed optimizations by adding them step by step on LLaMA-2-7B with a sequence length of 1024 and static eviction ratio of 75%. We make the following observations: 1) directly applying dynamic selection, e.g., LongCache to private LLM inference does not provide the expected efficiency improvement and even increases both latency and communication; 2) after static eviction, latency and communication of dynamic selection are reduced by 1.42$\times$ and 2.76$\times$, respectively; 3) our MPC-friendly optimizations, including clustering, linearization, reordering, and cross-layer index sharing further reduce the extra overhead introduced by dynamic selection without sacrificing the model performance; 4) MPCache eventually achieves 1.9$\times$ and 5.9$\times$ latency and communication reduction, respectively, and achieves better performance compared with H2O.
Figure 12: Effect of the number of adjacent layers.
Effect of hierarchical structures. To trade off the model performance and dynamic selection overhead, we evaluate different hierarchical structures on HotpotQA with a dynamic selection ratio of 20%. Specifically, we choose different cluster sizes $s$ and selection ratios at different levels (e.g., $s32(0.7)$ means selecting 70% clusters with $s=32$). From Table 6, we find that when the cluster size increases or the coarse-grained selection ratio decreases, the overhead becomes lower while the performance exhibits a downward trend. Moreover, appropriate course-grained selection may help improve the performance, e.g., the ratio changes from 90% to 50% with $s=32$;
Effect of the number of adjacent layers for cross-layer index sharing. In response to Section 3, we evaluate the efficiency-performance trade-off of the number of adjacent layers for cross-layer index sharing on HotpotQA in Figure 12. As observed, when the number of adjacent layers increases, the latency is reduced at the cost of the performance degradation.
Additional results. We present more experimental results, including the effect of $\alpha$, the necessity of KV cache, the discussion on 2PC protocol, and the comparison with average-based similarity approximation in Appendix F.
6 Conclusion
In this work, we propose an MPC-friendly KV cache eviction framework dubbed MPCache, that enables accurate and efficient private LLM inference. MPCache is a two-step framework combining static eviction and dynamic selection. To reduce the heavy overhead of dynamic selection, we propose a series of MPC-friendly optimizations. Extensive evaluations demonstrate that MPCache consistently outperforms prior-art KV cache eviction baselines across different generation tasks and significantly reduces both latency and communication overhead.
Appendix A Detailed Background and Related Works
A.1 Private LLM Inference
Recently, private LLM inference has attracted an increasing amount of research attention. Iron (Hao et al., 2022) uses coefficient encoding to compute homomorphic encryption (HE)-based convolution layers efficiently and uses MPC to compute non-linear layers. PUMA (Dong et al., 2023) proposes a series of 3PC protocols for both linear and non-linear functions to support private LLM inference, even under the scale of LLaMA-7B. BumbleBee (Lu et al., 2023) proposes HE-based protocols that enable the multiplication of large matrices and efficient protocols for non-linear functions similar to PUMA. CipherGPT (Hou et al., 2023) uses subfield vector oblivious linear evaluation (sVOLE) to reduce the communication of MatMuls significantly. BOLT (Pang et al., 2023) proposes a baby-step giant-step (BSGS) strategy that reduces the number of rotations on ciphertexts. SIGMA (Gupta et al., 2023) achieves private GPT inference with function secret sharing (FSS) and accelerates the computation on GPUs. PermLLM (Zheng et al., 2024) proposes an efficient protocol for non-linear functions based on the random permutation. However, they mainly focus on protocol optimization, and still incur significant overhead, especially on long sequences.
There are also works directly replacing expensive non-linear functions in Transformers, e.g., Softmax and GeLU with MPC-friendly operations. For instance, MPCFormer (Li et al., 2022) and Secformer (Luo et al., 2024) simplify Softmax by replacing exponential with an MPC-friendly quadratic function. MPCViT (Zeng et al., 2023) replaces exponential in Softmax with ReLU and selectively uses scaling attention using neural architecture search (NAS). SAL-ViT (Zhang et al., 2023) introduces external attention while PriViT (Dhyani et al., 2023) uses squaring function to replace Softmax. RNA-ViT (Chen et al., 2023) and Power-Softmax (Zimerman et al., 2024) use high-order polynomial approximation. Liu & Liu (2023) directly uniformly replaces GeLU with ReLU and replaces exponential in Softmax with ReLU. THE-X (Chen et al., 2022) approximates Softmax using an estimation model. Although these methods achieve high inference efficiency, they cannot avoid finetuning or re-training to preserve the model performance, making LLM development impractical. In conclusion, the above works still suffer from heavy overhead. Moreover, these works always handle full-length contexts during the LLM generation, incurring large latency and communication when handling long sequences. To solve this problem, our work aims to compress the KV cache in attention without training. Our proposed method can also be applied to different protocol frameworks for efficiency improvement.
A.2 KV Cache Compression
When tackling the LLM generation tasks, especially in long-context scenarios, the KV cache in the attention module becomes the most significant bottleneck due to the increasing sequence length. Therefore, how to effectively reduce the size of the KV cache is a high priority. System-level optimizations such as FlashAttention (Dao et al., 2022), FlashAttention-2 (Dao, 2023), FlashAttention-3 (Shah et al., 2024), and PagedAttention (Kwon et al., 2023) have been proposed to alleviate the problem. Meanwhile, many recent research efforts have been devoted to algorithm-level optimizations. For example, quantization methods (Hooper et al., 2024b; Zhang et al., 2024c; Kang et al., 2024; He et al., 2024; Liu et al., 2024d) have been proposed to compress KV cache to $1\sim 4$ bits, low-rank decomposition methods (Saxena et al., 2024; Chang et al., 2024; Zhao et al., 2024a; Zhang et al., 2024a) project the KV cache into low-rank space. In this work, we follow FlashAttention (Dao et al., 2022) to save the GPU memory and focus on another research line of algorithm-level optimization called KV cache eviction, which is designed to reduce the number of tokens and enable sparse attention without extra training.
KV cache eviction can be roughly categorized into 3 classes: 1) fixed-pattern algorithm: the position of important tokens is pre-defined before inference and remains consistent across decoding steps. However, this algorithm is not flexible for different LLMs and contexts (Xiao et al., 2023; Beltagy et al., 2020; Zaheer et al., 2020); 2) static algorithm: tokens are statically discarded and cannot be recovered in the subsequent decoding steps. This algorithm is usually efficient but suffers from significant performance degradation when the compression ratio is high (Zhang et al., 2024f; Liu et al., 2024c; Ge et al., 2023; Li et al., 2024b; Zhang et al., 2024d; Yang et al., 2024; Zhao et al., 2024b; Chen et al., 2024; Wan et al., 2024; Jiang et al., 2024; Oren et al., 2024); 3) dynamic algorithm: tokens are dynamically selected across different decoding steps. This algorithm is much more flexible but the dynamic selection usually involves more expensive operations (Xiao et al., 2024a; Liu et al., 2024b; Tang et al., 2024b; Chen et al., ). We quantitatively compare existing methods in Table 1.
Here, we introduce recent works of KV cache eviction. StreamingLLM (Xiao et al., 2023) proposes to keep a few initial tokens along with the recent tokens to recover the long-context performance. RazorAttention (Tang et al., 2024a) theoretically analyzes the scope of effective attention vision for each head. Scissorhands (Liu et al., 2024c), H2O (Zhang et al., 2024f), ALISA (Zhao et al., 2024b), spAtten (Wang et al., 2021), SnapKV (Li et al., 2024b) and LOOK-M (Wan et al., 2024), and TOVA (Oren et al., 2024) use the accumulated attention score of the historical tokens to select a small subset of KV cache. FastGen (Ge et al., 2023) and MInference 1.0 (Jiang et al., 2024) propose to allocate different eviction policies for different heads based on the sparsity pattern of the prompt. PyramidKV (Zhang et al., 2024d), PyramidInfer (Yang et al., 2024), and SqueezeAttention (Wang & Gan, 2024) consider allocating different KV cache budgets for different layers. InfLLM (Xiao et al., 2024a) and LongCache (Liu et al., 2024b) propose to dynamically select tokens based on the relationship between the current query and the key cache of previous tokens. RetrievalAttention (Liu et al., 2024a) establishes connections from the query to its nearest keys and the decoding query can first search its nearest query and then obtain the most relevant key vectors. LazyLLM (Fu et al., 2024) introduces an aux cache to enable selective KV cache eviction. Keyformer (Adnan et al., 2024) finds that the distribution after token pruning becomes uneven and proposes to smooth the distribution. The above works explicitly rely on the attention weights such that they are incompatible with FlashAttention (Luohe et al., 2024). To get rid of the dependency of attention weights, SirLLM (Yao et al., 2024) uses token entropy while Devoto et al. (2024) uses the L2-norm of the key cache to measure the token importance. However, these works are not designed or optimized for MPC since they either statically discard tokens that cause significant performance degradation, or dynamically select tokens that introduce more complex and MPC-unfriendly operations. Recently, learning-based methods (Xiao et al., 2024b; Gao et al., 2024) have also emerged, which require training and thus, are not the focus of this work.
A.3 Generative LLM Inference in Autoregressive-style
The generative inference procedure of LLM is generally in autoregressive style such as GPT-2 (Radford et al., 2019) and LLaMA (Touvron et al., 2023), and mainly consists of two stages: 1) the prefill (prompt) stage and 2) the decoding (generation) stage.
Prefill stage. The prefill stage serves as the first step of generation. LLM takes a prompt sequence as input and generates key-value (KV) caches for each layer. The attention can be computed as
$\mathbf{O}_{\texttt{prompt}}=\mathrm{Softmax}(\mathbf{Q}_{\texttt{prompt}}\cdot\mathbf{K}_{\texttt{prompt}}^{\top}/\sqrt{d})\cdot\mathbf{V}_{\texttt{prompt}},$ (3)
where $\mathbf{Q}_{\texttt{prompt}}\in\mathbb{R}^{H\times T\times d}$ denotes the query and $\mathbf{K}_{\texttt{prompt}}\in\mathbb{R}^{H\times T\times d},\mathbf{V}_{\texttt{prompt}}\in\mathbb{R}^{H\times T\times d}$ denote the key and value cache, respectively. After the prefill stage, the KV cache is generated as $\mathbf{K}_{\texttt{cache}}\leftarrow\mathbf{K}_{\texttt{prompt}}$ and $\mathbf{V}_{\texttt{cache}}\leftarrow\mathbf{V}_{\texttt{prompt}}$. KV cache retains previously computed key-value pairs, eliminating the need for costly re-computation of previous key and value vectors (Ott, 2019). Note that each layer is equipped with its unique KV cache and the generated KV cache is the foundation for the dowmstreaming decoding stage.
Decoding stage. The decoding stage uses and updates the stored KV cache to generate new tokens step-by-step. First, the KV cache is updated by concatenating new $\mathbf{k}\in\mathbb{R}^{H\times 1\times d}$ and $\mathbf{v}\in\mathbb{R}^{H\times 1\times d}$ as
$\mathbf{K}_{\texttt{cache}}\leftarrow\left[\mathbf{K}_{\texttt{cache}}||\mathbf{k}\right],~{}\mathbf{V}_{\texttt{cache}}\leftarrow\left[\mathbf{V}_{\texttt{cache}}||\mathbf{v}\right],$ (4)
[…]
$\mathbf{o}_{\texttt{dec}}=\mathrm{Softmax}(\mathbf{q}_{\texttt{dec}}\cdot\mathbf{K}_{\texttt{cache}}^{\top}/\sqrt{d})\cdot\mathbf{V}_{\texttt{cache}},$ (5)
where $\mathbf{q}_{\texttt{dec}}\in\mathbb{R}^{1\times d}$ denotes the current query. The attention output $\mathbf{o}_{\texttt{dec}}\in\mathbb{R}^{1\times d}$ is then sent to the multi-layer perceptron (MLP) layer for the subsequent computation.
Explanation about the use of KV cache. Although previous works overlook the use of KV cache (Dong et al., 2023; Lu et al., 2023, 2024), KV cache is already a fundamental component for existing generative LLMs to guarantee the inference correctness. If we do not use KV cache, the model has to re-compute all the previous tokens at each decoding step, significantly increasing the overhead. Hence, we apply KV cache to private LLM inference by default, and our aim is to reduce the computation associated with KV cache size for better efficiency. We also show the generation efficiency without KV cache in Section 5 and Appendix F.
Appendix B MPC Protocol Descriptions
B.1 Threat Model and Security
Consistent with previous works (Mohassel & Rindal, 2018; Li et al., 2022; Dong et al., 2023), MPCache adopts an honest-but-curious (a.k.a., semi-honest) security model in honest-majority (Lindell & Pinkas, 2009) where parties follow the protocol specifications but may also try to learn more from the information than allowed. In our threat model, we assume all the parties are aware of the LLM architecture and number of pruned tokens, which is consistent with HEPrune (Zhang et al., 2024e), Seesaw (Li et al., 2024a), SENet (Kundu et al., 2023), SNL (Cho et al., 2022), etc. We argue that this information does not compromise the client’s data or inference results, nor does it enable the client to access the model’s parameters.
B.2 2PC Protocol
We follow the 2PC protocols proposed in BumbleBee (Lu et al., 2023). The protocols are built based on the 2-out-of-2 additive secret sharing (SS), where secret value $x\in 2^{\ell}$ is shared by two random values $x_{0},x_{1}\in 2^{\ell}$ such that $x=x_{0}+x_{1}\pmod{2^{\ell}}$, and party $P_{i}$ gets $x_{i}$ (denoted as $[\![x]\!]$). SS supports both addition and multiplication on the secret shares. Without special declaration, we compute in $2^{\ell}$ and omit $\pmod{2^{\ell}}$ for brevity. In the case of $\ell>1$ (e.g., $\ell=64$) which support arithmetic operations (e.g., $+$, --, and $\cdot$), we refer to this type as arithmetic secret sharing. Boolean secret sharing refers to $\ell=1$ where (+,)(+,-)( + , - ) and $\cdot$ are replaced by bit-wise $\oplus$ and $\land$, respectively.
  • Addition. $[\![x+y]\!]$ can be computed as $(x_{0}+y_{0},x_{1}+y_{1})$, where $P_{i}$ can compute its share locally.
  • Multiplication. We write the multiplication of two shared values as $[\![xy]\!]=(x_{0}+x_{1})(y_{0}+y_{1})=x_{0}y_{0}+x_{1}y_{1}+x_{0}y_{1}+x_{1}y_{0}$ where two cross terms $x_{0}y_{1},x_{1}y_{0}$ can be computed using HE.
Lu et al. (2023) uses HE scheme that is based on ring learning-with-error (RLWE). For more details about the 2PC protocol, please refer to Lu et al. (2023); Ma et al. (2023).
B.3 3PC Protocol
We follow the 3PC protocols proposed in PUMA (Dong et al., 2023). The protocols are built based on the 2-out-of-3 replicated secret sharing (RSS), where a secret value $x\in 2^{\ell}$ is shared by three random values $x_{0},x_{1},x_{2}\in 2^{\ell}$ such that $x=x_{0}+x_{1}+x_{2}\pmod{2^{\ell}}$, and party $P_{i}$ gets $(x_{i},x_{i+1})$ (denoted as $[\![x]\!]$).
Let $(c_{1}$, $c_{2}$, $c_{3})$ be public constants, and $([\![x]\!],[\![y]\!])$ be two secret-shared values. The secure addition and multiplication procedures are as follows:
  • Addition. $[\![c_{1}x+c_{2}y+c_{3}]\!]$ can be computed as $(c_{1}x_{0}+c_{2}y_{0}+c_{3},c_{1}x_{1}+c_{2}y_{1},c_{1}x_{2}+c_{2}y_{2})$, where $P_{i}$ can compute its share locally. When $(c_{1}=1,c_{2}=1,c_{3}=0)$, we get $[\![x+y]\!]$.
  • Multiplication. Parties follow steps: i) first, $P_{i}$ computes $z_{i}=x_{i}y_{i}+x_{i+1}y_{i}+x_{i}y_{i+1}$ locally; ii) parties then perform re-sharing by letting $P_{i}$ sends $z_{i}^{\prime}=\alpha_{i}+z_{i}$ to $P_{i-1}$, where $\alpha_{0}+\alpha_{1}+\alpha_{2}=0$ ($P_{i}$ can generate $\alpha_{i}$ using pseudorandom generators with negligible overhead as Mohassel & Rindal (2018)); iii) finally, $\{(z_{0}^{\prime},z_{1}^{\prime}),(z_{1}^{\prime},z_{2}^{\prime}),(z_{2}^{\prime},z_{0}^{\prime})\}$ form the 2-out-of-3 replicated secret shares of $[\![xy]\!]$.
For more details about the 3PC protocol, please refer to Mohassel & Rindal (2018); Ma et al. (2023).
B.4 Token Gathering Protocol
Token gathering is used to retrieve tokens in the KV cache based on the indices, which has the same functionality as $\mathrm{torch.gather(tensor,indices)}$ in PyTorch programming. We illustrate the overall procedure in Figure 13.
For brevity, in Algorithm 2, we show the pipeline that retrieves one token from the key cache (we also omit the head dimension for simplification). The first step is converting a ciphertext index $[\![\mathbf{id}]\!]$ into a ciphertext one-hot vector $[\![\mathbf{o}]\!]\in\mathbb{R}^{1\times T}$ based on equal protocol $\Pi_{\mathrm{Equal}}$, where $T$ denotes the number of tokens. Given a ciphertext key cache $[\![\mathbf{K}]\!]\in\mathbb{R}^{T\times d}$, where $d$ denotes the embedding dimension, multiplying $[\![\mathbf{o}]\!]$ with $[\![\mathbf{K}]\!]$ (matrix-vector multiplication $\Pi_{\mathrm{MatVec}}$) can generate an output with a dimension of $1\times d$, which is the retrieved token. To extend the case to retrieve $m$ tokens, we concatenate $m$ one-hot vectors to form a matrix $[\![\mathbf{O}]\!]\in\mathbb{R}^{m\times T}$, and then multiply $[\![\mathbf{O}]\!]$ with $[\![\mathbf{K}]\!]$ (matrix-matrix multiplication $\Pi_{\mathrm{MatMul}}$) to generate an output with a dimension of $m\times d$, which is the retrieved $m$ tokens.
Note that token gathering protocol is also used on the value cache, and its indices are consistent with that of the key cache.
Figure 13: Illustration of token gathering procedure where we give an example of retrieving 3 tokens from 6 tokens.
Algorithm 2 Token gathering protocol $\Pi_{\mathrm{Gather}}$ for retrieving one token
Input :
A ciphertext key cache $[\![\mathbf{K}]\!]\in\mathbb{R}^{T\times d}$ and a ciphertext index $[\![\mathbf{id}]\!]$.
Output :
Key cache $[\![\mathbf{K}]\!]^{\prime}\in\mathbb{R}^{1\times d}$ with the selected token.
for $i\in[0,\dots,T-1]$ do
2       Parties jointly generate the one-hot vector as $[\![\mathbf{o}[i]]\!]=\Pi_{\mathrm{Equal}}([\![\mathbf{id}]\!],i)$;
Parties jointly compute the retrieved key cache as $[\![\mathbf{K}]\!]^{\prime}=\Pi_{\mathrm{MatVec}}([\![\mathbf{o}]\!],[\![\mathbf{K}]\!])$; return $[\![\mathbf{K}]\!]^{\prime}$.
B.5 Paralleled Top-k Protocol
We follow a similar top-k protocol as CipherGPT (Hou et al., 2023). However, the top-k protocol in CipherGPT is designed for selecting k elements from a vector and not optimized with parallelization, and hence, for KV cache selection in this work, multiple attention heads need to be computed sequentially, leading to significant overhead. Therefore, we improve the top-k protocol with parallelization along the attention head axis to support the head-wise selection, which is more efficient than the original top-k protocol in CipherGPT.
Appendix C Observation from Pattern Discovery of Large Attention Maps
It is sufficient to use a few tokens within the observation window to distinguish the attention patterns since the structure of attention maps is stable at different decoding steps (Liu et al., 2024c; Yang et al., 2024; Ge et al., 2023; Li et al., 2024b). In Figure 14, we visualize the large attention map with hundreds of tokens on the PiQA (Bisk et al., 2020) dataset to further verify our observation in Section 3. As can be observed, there are three types of tokens defined in Section 3: 1) IA tokens in red blocks which usually appear as attention sinks mentioned in StreamingLLM (Xiao et al., 2023). 2) IC tokens in orange blocks; 3) UIA tokens in blue blocks; The pattern of attention maps motivates us to statically discard the UIA tokens which may have negligible impact on further generation, and dynamically select important tokens from IC tokens at each decoding step for sparse attention computation.
Figure 14: Attention patterns across different layers on LLaMA-7B.
Appendix D Pseudocode of MPCache Framework
We describe the overall flow of our MPCache in detail as shown in Algorithm 3.
Algorithm 3 MPCache: KV cache eviction combining static and dynamic algorithm
Input :
Input sequence prompt; LLM model $\mathcal{M}$; number of layers and attention heads $L$ and $H$; dynamic selection ration $\eta\in[0,1]$; three types of tokens IA, IC, and UIA (introduced in Section 3); cluster size $s$; decoding steps $E$.
Output :
Attention output and evicted KV cache.
Step 1: Look-once static eviction during prefill stage
for $l\in[0,\dots,L-1]$ do
4       $\mathbf{Q}^{(l)},\mathbf{K}^{(l)},\mathbf{V}^{(l)}\leftarrow\mathcal{M}(\texttt{prompt})$; if $l\%2==0$ then
5             $\mathbf{ATTN}_{\texttt{window}}^{(l)}\leftarrow\mathrm{Softmax}(\mathbf{Q}^{(l)}[:,-len(\texttt{prompt})\times 0.2:,:]\cdot\mathbf{K}^{(l)\top})$; $\{IA,IC,UIA\}^{(l)}\leftarrow\mathrm{static\_evict}(\mathbf{ATTN}_{\texttt{window}}^{(l)})$;
6      else
7             $\{IA,IC,UIA\}^{(l)}\leftarrow\{IA,IC,UIA\}^{(l-1)}$; $\triangleright$ Cross-layer index sharing (Section 4.4)
8      $\mathbf{K}^{(l)}\leftarrow\mathrm{token\_gather}(\mathbf{K}^{(l)},index=\{IA,IC\}^{(l)})$; $\mathbf{V}^{(l)}\leftarrow\mathrm{token\_gather}(\mathbf{V}^{(l)},index=\{IA,IC\}^{(l)})$;
9Step 2: Query-aware dynamic selection during decoding stage for $e\in[0,\dots,E-1]$ do
10       for $l\in[0,\dots,L-1]$ do
11             if $l\%2==0$ then
12                   $\mathbf{sim}\leftarrow\mathrm{Sim}(\mathbf{q}^{(l,e)},\mathbf{K}^{(l,e)},cluster\_{size}=s)$; $\triangleright$ Follow Equation (2) $\mathbf{index}^{(l,e)}\leftarrow\mathrm{topk}(\mathbf{sim},k=\mathbf{K}^{(l,e)}.size()\times\eta)$;
13            else
14                   $\mathbf{index}^{(l,e)}\leftarrow\mathbf{index}^{(l-1,e)}$; $\triangleright$ Cross-layer index sharing (Section 4.4)
15            $\mathbf{K}^{(l,e)}\leftarrow\mathrm{token\_gather}(\mathbf{K}^{(l,e)},index=\mathbf{index}^{(l,e)})$; $\mathbf{V}^{(l,e)}\leftarrow\mathrm{token\_gather}(\mathbf{V}^{(l,e)},index=\mathbf{index}^{(l,e)})$; $\mathbf{O}^{(l,e)}\leftarrow\mathrm{Softmax}(\mathbf{q}^{(l,e)}\cdot\mathbf{K}^{(l,e)\top}/\sqrt{d})\cdot\mathbf{V}^{(l,e)}$; $\triangleright$ Sparse attention
16
return $\mathbf{O,K,V}$.
Appendix E Paradigm Comparison with Dynamic Algorithm
We compare the paradigm between the dynamic algorithm and our MPCache as shown in Figure 15. Through performing static eviction during the prefill stage, we improve the efficiency of all decoding steps since the token selector (i.e., similarity approximation and top-$k$ ranking) only needs to handle a smaller number of tokens during the decoding stage (30% in this case).
Figure 15: Paradigm comparison between (a) dynamic algorithm and (b) our proposed MPCache combining static eviction and dynamic selection. MPCache discards unimportant tokens to reduce the decoding overhead (red texts mean the differences).
Appendix F Supplemental Experiments
F.1 Supplemental Setups
Experimental environment. The model performance is evaluated with LongBench on an NVIDIA A100 80GB GPU in PyTorch. The latency is evaluated with Secretflow under the LAN setup (Rathee et al., 2020) with 377MBps bandwidth and 0.3ms echo latency (Rathee et al., 2020) on Intel(R) Xeon(R) Gold 5220R CPU @ 2.20GHz.
KV cache clustering configuration. For hierarchy, we in practice choose a two-level hierarchical structure, i.e., $n=2$, and when the final dynamic selection ratio $\alpha<0.5$, we drop 50% clusters at the 1st hierarchical level. For the XSUM dataset, we use a cluster size of 8 at the 1st hierarchical level and 4 at the 2nd hierarchical level. For the long-context LongBench, we use larger clusters, i.e., 32 at the 1st hierarchical level and 16 at the 2nd hierarchical level.
Static eviction. During the static eviction, we only compute the attention using the last 20% tokens in the prompt. However, using 20% tokens still incurs a CUDA out-of-memory (OOM) error when processing a long-context prompt (i.e., longer than 24k tokens). To solve this problem, we adaptively adjust the ratio to 10% tokens instead. We also notice that the choice of the ratio won’t cause significant performance fluctuations, so we omit the discussion in this paper.
How to determine the hyper-parameters, e.g., eviction ratios in the realistic scenario? In practice, we believe the static and dynamic eviction ratio should be determined based on two aspects: 1) the expected sequence length and 2) the latency that can be tolerated/afforded by the server and the client. Our framework proposes a novel algorithm with configurable hyper-parameters to enable exploring the Pareto front of the LLM performance and efficiency. The framework has been validated across different benchmarks and different LLMs, which demonstrates its value and potential for practical use cases.
F.2 Model Architectures
We list the model architectures used in this work in Table 7, including GPT-2444https://huggingface.co/openai-community/gpt2, LongChat-V1.5-7B [^5] https://huggingface.co/lmsys/longchat-7b-v1.5-32k, and LLaMA-3-8B-Instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. LLaMA-3 is equipped with grouped-query attention (GQA) (Ainslie et al., 2023).
Table 7: LLM architectures used in this work.
Model # Layers # Attention Heads # KV Heads Embedding Dimension Max Length Non-linear Function # Parameters
GPT-2 12 12 12 768 1024 GeLU 124M
LongChat-V1.5-7B 32 32 32 4096 32K SiLU 7B
LLaMA-3-8B-Instruct 32 32 8 4096 8K SiLU 8B
F.3 Detailed Descriptions about Baselines
StreamingLLM (Xiao et al., 2023) follows a fixed eviction pattern, i.e., keep the local tokens and initial tokens) across different decoding steps. H2O (Zhang et al., 2024f) and TOVA (Oren et al., 2024) statically prune the KV cache and these discarded tokens cannot be recovered at subsequent decoding steps. SnapKV (Li et al., 2024b) statically prunes the KV cache only during the prefill stage. InfLLM (Xiao et al., 2024a) employs block-level dynamic token selection during the decoding stage. It requires selecting several representative tokens within a block and computing the relevance score using these representative tokens. LongCache (Liu et al., 2024b) uses the idea of cosine similarity between the query and key cache of previous tokens to select relevant tokens without static eviction and token clustering. Note that LongCache separates the positional embedding (PE) from the KV cache. Since our focus is on the dynamic selection metric in this work, we do not apply PE separation in LongCache.
F.4 Necessity of KV Cache in MPC-based LLM
KV cache plays a crucial role in storing context information that the model deems relevant for subsequent token generation. KV cache eliminates the need for costly re-computation of previous key and value vectors (Ott, 2019). While the existing MPC framework PUMA (Dong et al., 2023) builds protocols to support private LLM inference, it still requires more than 40 seconds and 800 MB to generate one token with a sequence length of 6 on GPT-2. The weakness is that it does not use KV cache, so it requires to re-compute the entire sequence for each token decoding. As shown in Table 8, KV cache for LLM generation is essential in private inference, bringing significant efficiency improvement, especially for longer sequences.
Table 8: Per-token generation efficiency on GPT-2 with and without KV cache.
Method Sequence Length=6 Sequence Length=16
Latency (s) Communication (GB) Latency (s) Communication (GB)
w/o KV cache 19.78 0.273 44.40 0.872
w/ KV cache 7.890 0.068 8.313 0.071
F.5 Supplemental Ablation Studies
Effect of hyper-parameter $\alpha$. To study how $\alpha$ impacts the similarity approximation, we select different $\alpha$’s on different datasets as shown in Figure 17. As can be observed, although the effects of different $\alpha$ do not occur in a certain pattern, we can still discover some patterns related to the dataset from the trend in Figure 17: on TriviaQA, the model may prefer larger $\alpha$ while it may prefer smaller $\alpha$ on HotpotQA instead. Since $\alpha=0.6$ shows relatively better performance in these cases, we choose $\alpha=0.6$ by default in our experiments.
Figure 16: The influence trend of similarity approximation with different $\alpha$ values ranging from 0 to 1.
Figure 17: Comparison with average-based similarity approximation. MP means maximum dot product.
Figure 18: Extension MPCache to 2PC protocol.
Comparison with average-based similarity approximation. A straightforward and efficient way to aggregate the information of a key cache cluster is the average. We compare our proposed method ($\alpha=0.6$) with average-based similarity on the XSUM dataset with a cluster size of 16 in Figure 17. Specifically, we perform dynamic selection with different ratios after 75% tokens are statically discarded. As can be observed, using average suffers from significant performance degradation under different ratios. With the compression ratio increasing, the degradation of the average-based method becomes more serious. An intuitive explanation is that using the average of a cluster may make some important tokens averaged and ignored. In contrast, our approximation can effectively maintain the model performance. We theoretically analyze the similarity approximation algorithm in Appendix G.
Discussion on 2PC protocol. We evaluate the 2PC efficiency in Figure 18. It is observed that MPCache achieves 1.63×1.63$\times$1.63 × and 1.79×1.79$\times$1.79 × latency and communication reduction compared with the full cache, and 2.58$\times$ and 2.48×2.48$\times$2.48 × latency and communication reduction compared with LongCache. Since the multiplication communication in 2PC is larger than in 3PC, the cost of similarity approximation becomes higher. To solve this problem, we can leverage the random projection based on Johnson-Lindenstrauss (JL) Lemma (Johnson et al., 1986) to reduce the dimensionality while preserving the token distance. We leave the research as our future work.
Appendix G Theoretical Analysis of Similarity Approximation
As mentioned in Section 2, given query $\mathbf{q}\in\mathbb{R}^{H\times 1\times d}$, key cache $\mathbf{K}\in\mathbb{R}^{H\times T\times d}$, and value cache $\mathbf{V}\in\mathbb{R}^{H\times T\times d}$, the overall goal of KV cache eviction is to find an optimal policy $\mathcal{P}$ to minimize the gap between the attention outputs (here we omit $\mathbf{V}$ for simplification) as
$\mathcal{P}^{*}=\mathrm{argmin}|\mathrm{Softmax}(\mathbf{q}\cdot\mathbf{K}^{\top})-\mathrm{Softmax}(\mathbf{q}\cdot\mathbf{K}^{\prime\top})|,$ (6)
where $\mathbf{K}^{\prime}$ is a subset of $\mathbf{K}$ selected by $\mathcal{P}$. However, when grouping $\mathbf{K}$ into clusters for efficiency, the problem becomes challenging. We denote the key cache cluster as $\mathbf{K}_{c}$ (cluster size is $s$), and our goal is to find a way to accurately approximate the similarity between $\mathbf{q}$ and the key cluster $\mathbf{K}_{c}$. This problem is equivalent to “how can we effectively aggregate the cluster information to obtain a cluster representation and measure its importance?”
We assume there exists a function $\phi$ that aggregates the key cluster $\mathbf{K}_{c}$, and we define the optimization problem as
$\min|\sum_{j=0}^{s-1}\exp(\mathbf{q}\cdot\mathbf{K}_{cj}^{\top})-\mathbf{q}\cdot\phi(\mathbf{K}_{c}^{\top})|.$ (7)
Drawback of average-based clustering. As mentioned, the simplest way to represent the cluster is the average and $\phi(\mathbf{K}_{cj}^{\top})$ becomes $\sum_{j=0}^{s-1}\mathbf{K}_{cj}^{\top}/s$. This happens to be equivalent to directly drop $\exp$ in $\sum_{j=0}^{s-1}\exp(\mathbf{q}\cdot\mathbf{K}_{cj}^{\top})$, introducing information loss. Intuitively, if there are tokens with low importance and tokens with high importance within one cluster, the overall importance will be averaged, leading to the neglect of the crucial tokens.
Different from the average-based method, using the max dot product aims to protect the crucial tokens with large scores as much as possible. This observation is aligned with Tang et al. (2024b).
$\mathrm{MaxDotProduct}:~{}~{}\mathbf{q}\cdot\phi(\mathbf{K}_{c}^{\top})=\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}\cdot\mathbf{k}.$ (8)
In order to approximate $\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}\cdot\mathbf{k}$ without accessing all the tokens in $\mathbf{K}_{c}$, we utilize the bounding volume proposed by Klosowski et al. (1998). Our optimizations are described in Section 4.3.
Appendix H Overall Secure Inference Framework
Take 2PC as an example in Figure 19, we illustrate the secure inference framework following Lu et al. (2023) where the server owns the proprietary LLM parameter and the client possesses private input data. During inference, the data is secretly shared between two parties. Linear layers are computed using the HE protocol, and non-linear layers require interactive protocols between the two parties based on oblivious transfer (OT) and HE. Figure 19 illustrates the detailed data flow of MPCache during the token generation.
Figure 19: Overall secure inference data flow of MPCache during token generation.