### v2

# For Efficient Private LLM Inference

Abstract
Private large language model (LLM) inference based on secure multi-party computation (MPC) achieves formal data privacy protection but suffers from significant latency overhead, especially for long input sequences. While key-value (KV) cache eviction and sparse attention algorithms have been proposed for efficient LLM inference in plaintext, they are not designed for MPC and cannot benefit private LLM inference directly. In this paper, we propose an accurate and MPC-friendly KV cache eviction framework, dubbed MPCache, building on the observation that historical tokens in a long sequence may have different effects on the downstream decoding. Hence, MPCache combines a look-once static eviction algorithm to discard unimportant KV cache and a query-aware dynamic selection algorithm to activate only a small subset of KV cache for attention computation. MPCache further incorporates a series of optimizations for efficient dynamic KV cache selection, including MPC-friendly similarity approximation, hierarchical KV cache clustering, and cross-layer index-sharing strategy. Extensive experiments demonstrate that MPCache consistently outperforms prior-art KV cache eviction baselines across different generation tasks and achieves $1.8\sim 2.01\times$ and $3.39\sim 8.37\times$ decoding latency and communication reduction on different sequence lengths, respectively. The code can be found here.
1 Introduction
Large language models (LLMs) have recently demonstrated remarkable ability in a wide range of applications such as document summarization (huang2021efficient, ; narayan2018don, ; zhang2024benchmarking, ), question answering (kovcisky2018narrativeqa, ; dasigi2021dataset, ; yang2018hotpotqa, ), and dialogue systems (thoppilan2022lamda, ; chiang2023vicuna, ; taori2023stanford, ). However, LLM-based machine learning as a service (MLaaS) on the cloud has raised serious privacy concerns as the users are required to upload their prompts to the cloud, which may contain sensitive personal information. Meanwhile, the service provider is unwilling to offload the trained model to the user to protect the proprietary model weights. In recent years, secure multi-party computation (MPC)-based LLM inference has become a prevailing and mainstream research trend to address privacy concerns lu2023bumblebee ; hou2023ciphergpt ; dong2023puma ; gupta2023sigma ; pang2023bolt ; rathee2024mpc ; zheng2024permllm ; liu2023llms , which enables the users and the cloud to conduct the LLM inference jointly, but nothing else can be derived beyond the final inference results.
However, MPC-based LLM inference illustrated in Figure 1(a) faces serious efficiency challenges, especially for long sequences. We profile the decoding efficiency of GPT-2 with the Secretflow framework (ma2023secretflow, ) using recent 2-party computation (2PC) (lu2023bumblebee, ) and 3-party computation (3PC) protocols (dong2023puma, ). As can be observed in Figure 1(b) and (c), attention computation dominates the latency and communication for both 2PC and 3PC protocols. Moreover, Softmax accounts for the majority of the overall cost due to its complex operators including max, exponential, and division, especially with an increasing sequence length.
Figure 1: (a) MPC-based LLM inference. (b) Breakdown of decoding latency and communication with a sequence length of 512. Attention dominates the overhead for both 3PC and 2PC protocols. (c) The cost of Softmax scales with the sequence length rapidly. (d) Inference cost before and after KV cache eviction. Blocks in slash indicate the extra overhead introduced by eviction.
To reduce the cost of private LLM inference, previous works focus on developing more efficient MPC protocols (lu2023bumblebee, ; dong2023puma, ; pang2023bolt, ; hou2023ciphergpt, ), replacing non-linear activation functions with more MPC-friendly operators (liu2023llms, ; li2022mpcformer, ; zeng2023mpcvit, ), or directly modifying the model architecture (rathee2024mpc, ). However, they still incur significant overhead or require expensive finetuning or re-training, and cannot be directly applied to LLMs. Another line of works leverages key-value (KV) cache eviction (sparse attention) to reduce the number of tokens involved in the attention computation (zhang2024h2o, ; ge2023model, ; liu2024scissorhands, ; zhao2024alisa, ; zhang2024pyramidkv, ; fu2024lazyllmdynamictokenpruning, ). Although they have demonstrated significant memory and computation reduction for plaintext LLM inference without the need of finetuning, they are not friendly to MPC. As shown in Figure 1(d), directly applying the dynamic KV cache eviction algorithm (liu2024farewell, ) incurs even more communication and latency overhead over the baseline model since it introduces expensive operators in MPC, including top-$k$ ranking, token gathering, etc, as elaborated in Section 3. Therefore, there is an urgent need for a training-free MPC-friendly algorithm designed for private LLM inference.
To overcome the heavy overhead of attention computation, we make the following observations that motivate our MPCache: 1) LLM attention maps are overall sparse for long input prompts, motivating us to perform static eviction and directly prune the KV cache of unimportant tokens; 2) LLM attention shows token-wise locality (liu2023cachegen, ), motivating us to build an efficient clustering algorithm for dynamic selection of the KV cache; 3) LLM attention of adjacent layers shows similar patterns, motivating us to share the KV cache selection for adjacent layers to further improve efficiency. Our contributions can be summarized as follows:
  • We observe the cost of MPC-based LLM inference mainly comes from attention computation and propose MPCache, an MPC-friendly KV cache eviction framework to reduce the LLM inference latency and communication.
  • We identify the challenges when applying KV cache eviction to MPC. To tackle the problems, MPCache combines look-once static eviction and query-aware dynamic selection with a series of optimizations, including MPC-friendly similarity approximation, hierarchical KV cache clustering, and cross-layer index-sharing strategy.
  • With extensive experiments, we demonstrate the performance of MPCache consistently exceeds the prior-art KV cache eviction algorithms across different generation tasks and achieves up to $2.01\times$ and $8.37\times$ decoding latency and communication, respectively.
2 Problem Formulation and Background
2.1 Problem Formulation
Generative LLM inference can be divided into prefill and decoding stages (refer to Appendix A.3). We formally describe the decoding process with KV cache eviction in Algorithm 1. The KV cache eviction policy, denoted as $\mathcal{P}$, aims to minimize the attention computation by only preserving a subset of tokens, which typically involves three steps: 1) $\mathcal{P}$ first computes the similarity between the query and key cache of previous tokens (line # 1); 2) $\mathcal{P}$ then ranks the previous tokens based on the similarity score and applies the top-$k$ algorithm to determine the indices of relevant tokens (line # 2); 3) the KV cache is then retrieved based on the indices, denoted as token gathering (line # 3) [^1] We describe the procedure and protocol of token gathering in Appendix B.4., followed by sparse attention computation with the selected KV cache (line # 4). To compute the similarity in line # 1, existing works have used accumulated attention score of the historical tokens (liu2024scissorhands, ; zhang2024h2o, ; zhao2024alisa, ; yang2024pyramidinfer, ; zhang2024pyramidkv, ) or cosine similarity (liu2024farewell, ; xiao2024infllm, ).
1
2
Input :  Query, key, and value cache $\mathbf{q},\mathbf{K,V}$.
Output : Sparse attention output $\mathbf{O}$
3
41ex
[…]
9return $\mathbf{O}$.
Algorithm 1 Formulation of KV cache eviction
KV cache eviction reduces the attention computation complexity from $\mathcal{O}(Td)$ to $\mathcal{O}(kd)$, where $T,d$ denote the sequence length and embedding dimension, respectively, and $k\ll T$. However, it introduces MPC-unfriendly operations, including similarity approximation, top-$k$ ranking, and token gathering, hindering its benefits in MPC-based LLM inference. Hence, the goal of this work can be summarized as follows: “How can we design an MPC-friendly KV cache eviction algorithm $\mathcal{P}$ to minimize MPC-based LLM inference overhead without sacrificing LLM performance?”
Figure 2: (Upper) token types in attention maps where means the token is selected and means the token is not selected. (Lower) three types can be observed in attention map with more tokens.
2.2 Background
KV cache eviction. There has been a surge in improving the efficiency of private LLM inference. Existing works focus on the protocol optimization (pang2023bolt, ; dong2023puma, ; lu2023bumblebee, ) or replacing non-linear functions with MPC-friendly operators (li2022mpcformer, ; zeng2023mpcvit, ; dhyani2023privit, ). However, they either incur large overhead for long sequences or require expensive training. KV cache eviction has been widely explored for plaintext inference and can be mainly classified into 3 categories: 1) fixed-pattern algorithms like xiao2023efficient ; beltagy2020longformer always keep the tokens in the same position, e.g., initial and recent tokens across decoding steps, lacking flexibility for different LLMs and contexts; 2) static algorithms like zhang2024h2o ; li2024snapkv ; ge2023model discard tokens based on the accumulated attention scores of historical tokens, which are efficient but suffers from large performance degradation when the compression ratio is high; 3) dynamic algorithms like xiao2024infllm ; tang2024quest ; hooper2024squeezed compute the similarity between the current query and key cache for every decoding step, which is more accurate but requires repetitive token selection at each step. Recently, there are learning-based methods xiao2024duoattention ; gao2024seerattention which rely on the training process. Different from prior works in Table 1, MPCache is a training-free and hybrid method and equipped with MPC-friendly optimizations, simultaneously achieving high efficiency and performance. We leave a detailed review of existing works in Appendix A.
Table 1: Qualitative comparison with prior works of attention optimization.
Representative
Work
Category
Token Similarity
Approximation
Top-$k$
Ranking
Token
Gathering
Layer-wise
Optimization
MPC
Efficiency
Model
Performance
li2022mpcformer ; zeng2023mpcvit ; zhang2023sal
Softmax
Approximation
Training
Required
Impractical
for LLM
xiao2024duoattention ; yuan2025native ; yang2025lserve Learning-based - - - -
Training
Required
High
xiao2023efficient ; beltagy2020longformer ; han2023lm Fixed-pattern - - - - High Low
li2024snapkv ; wan2024look ; cai2024pyramidkv Static Accumulated Attn. Usually Once Token-wise - High Low
liu2024farewell ; tang2024quest ; chenarkvale Dynamic Token-wise Similarity Token-wise & Step-wise Token-wise - Low High
MPCache (ours)
Static+Dynamic
Hierarchical Clustering
Cluster-wise Similarity
Cluster-wise & Step-wise
Cluster-wise
Cross-layer
Index-sharing
High High
MPC and threat model. MPC (goldreich1998secure, ) is a cryptographic technique recently developed to enable LLM inference while protecting the privacy of both data and model, as shown in Figure 1(a). In an MPC framework, to protect a certain tensor, it is often split into multiple secret shares and distributed across different parties involved in the computation (lu2023bumblebee, ; dong2023puma, ; mohassel2018aby3, ; hou2023ciphergpt, ). We adopt an honest-but-curious threat model and apply MPCache to both 2PC and 3PC protocols, which involve 2 parties and 3 parties in the computation, respectively. We refer interested readers to Appendix B, where the threat model and underlying protocols are more clearly explained.
3 Motivations and Challenges
In this section, we discuss the key observations that motivate the design of MPCache.
Figure 3: Motivating inspirations. (a) Statically evicting 60% tokens during the prefill stage still maintains the performance; (b) less than 20% tokens contribute to decoding; (c) cross-layer commonality among different numbers of adjacent layers.
Observation 1: the attention map of a long input sequence is usually sparse, and the KV cache of historical tokens demonstrates different impacts over the downstream decoding. We show the attention map of different heads and layers of LLaMA-2-7B in Figure 2 and leave visualizations of larger attention maps in Appendix C. From Figure 2, we can classify different tokens into 3 categories: 1) important to all tokens (IA in red box): the attention scores remain high for the entire column, e.g., 0th and 1st columns in Figure 2(a), indicating these tokens are important for the generation of all downstream tokens and hence, need to be always preserved; 2) un-important to all tokens (UIA in blue box): the attention scores remain low for the entire column, e.g., 2nd and 3rd columns in Figure 2(a), indicating these tokens can be discarded without impacting the downstream decoding; 3) important to certain tokens (IC in orange box): the attention scores vary for different tokens, e.g., 4th and 5th columns in Figure 2(a), indicating these tokens impact a subset of downstream tokens, and hence, cannot be directly pruned.
We verify the observations on LLaMA-2-7B with different input sequence lengths. As shown in Figure 3(a), almost 60% tokens can be statically evicted while preserving the LLM performance. While further pruning the remaining tokens starts to degrade the LLM performance, as shown in Figure 3(b), in each decoding step, only less than 20% of the remaining tokens contribute to the decoding. The above observation motivates us to statically evict the KV cache of UIA tokens and dynamically select a subset of IC tokens at each decoding step.
Observation 2: dynamic KV cache selection incurs non-negligible overhead in MPC. While dynamic KV cache selection is accurate and reduces the attention computation cost, it incurs non-negligible overhead due to MPC-unfriendly operations. In Figure 1(d), we show the extra overhead when 5% tokens are dynamically selected. The MPC-unfriendly operations mainly include:
  • Similarity computation (Algorithm 1 line # 1): cosine similarity is widely used, which requires computing the multiplication between the current query with the key cache of all previous tokens;
  • Top-$k$ ranking (Algorithm 1 line # 2): to compute the indices of relevant tokens, top-$k$ is usually inevitable (zhang2024h2o, ; ge2023model, ; zhao2024alisa, ). Unlike plaintext inference, top-$k$ ranking in MPC involves frequent comparison protocol, which incurs high latency and communication cost (rathee2020cryptflow2, ).
  • Token gathering (Algorithm 1 line # 3): after the top-$k$ ranking, the KV cache of selected tokens is gathered based on the indices. Unlike plaintext inference, such gathering protocol in MPC is much more expensive since both KV cache and indices are ciphertexts. Therefore, as described in Algorithm 2, each index is first converted to a one-hot vector and then multiplied with the KV cache, requiring repetitive MPC-unfriendly equal protocols.
Inspired by token-wise locality (liu2023cachegen, ; zhu2023biformer, ), our key insight is to group the adjacent tokens into clusters, which can reduce the complexity of dynamic selection in proportion to the cluster size. However, this introduces extra questions on how to measure the similarity between a cluster and the current query, how to build the cluster, etc, which is discussed in Section 4.3.
Figure 4: Overview of our proposed MPCache.
Observation 3: adjacent layers share similar top-$k$ ranking of KV cache, providing an extra opportunity for efficiency optimization. Due to the residual, we hypothesize adjacent layers may share a similar top-$k$ ranking of the KV cache. To verify the assumption, we define commonality score to measure the ratio of common top-$k$ indices of $m$ adjacent layers as
$\frac{1}{k(L-m)}\sum_{l=1}^{L-m}\left|\bigcap_{i=l}^{l+m}\mathbf{idx}_{i}[:k]\right|,$
where $\mathbf{idx}_{i}[:k]$ denotes the set of top-$k$ indices for $i$-th layer, $L$ is the number of layers, and $\left|\cdot\right|$ counts the number of elements in a set. As shown in Figure 3(c), adjacent layers demonstrate a high similarity of top-$k$ indices, which indicates the query tends to focus on the KV cache of the similar tokens. The score reduces when $m$ increases, which motivates us to share the indices of selected tokens among adjacent layers to trade off efficiency and performance.
4 MPCache: MPC-friendly LLM Inference
4.1 Overview of MPCache
Overall design. Driven by the observations, we propose an MPC-friendly KV cache eviction framework, dubbed MPCache  as shown in Figure 4. It consists of two steps: 1) look-once static eviction during the prefill stage to discard the UIA tokens (Section 4.2); 2) query-aware dynamic selection during the decoding stage to select only a small subset of the remaining IC tokens for sparse attention (Section 4.3). A series of MPC-friendly optimizations are proposed to reduce the overhead of selection (Section 4.3 and 4.4). The pseudocode is shown in Algorithm 3 and the overall data flow is shown in Appendix H.
Symbol definition. For clarity, we summarize the symbols used in this paper. We define $L$ as the number of layers, $H$ as the number of attention heads, $T$ as the number of tokens, $d$ as the embedding dimension, $s$ as the cluster size, and $C$ as the number of clusters.
4.2 Step 1: Look-once Static KV Cache Eviction
Figure 5: The illustration of static eviction.
To prune the KV cache of UIA tokens as observed in Section 3, we use static eviction once during the prefill stage. To measure the token importance and identify UIA tokens, we compute the attention map and then, accumulate the attention scores for each token. Similar to li2024snapkv ; zhao2024alisa , we find it is sufficient to only sum up the scores of the last 20% tokens in the prompt to compute the accumulated attention score. Then, we rank the accumulated attention scores to select the important tokens with the highest scores and discard the UIA tokens.
Protocol complexity analysis. Compared to the baseline computation of the prefill stage, static eviction only involves accumulating the attention scores, which are locally computed without any communication, and top-$k$ ranking. Because the static eviction is performed only once, the cost can be amortized by the entire generation process, and hence, becomes negligible. Meanwhile, with UIA tokens pruned, the efficiency of the dynamic selection can be improved for all decoding steps and overall efficiency.
4.3 Step 2: MPC-friendly Dynamic KV Cache Selection
Figure 6: Hierarchical and dynamic KV cache clustering and selection procedure.
To reduce the overhead of dynamic token selection as shown in Figure 1(c), we propose to group the KV cache of adjacent tokens into clusters in Figure 6(b). The most important question is “how to aggregate the information of a cluster and measure the cluster importance accurately and efficiently?”
MPC-friendly similarity approximation with clustering. A naive method for similarity approximation is to compute the average of the key cache within a cluster and directly compute the cosine similarity with the average. However, as shown in Figure 7, the naive approach incurs large performance degradation. Our intuition is that the approximation should preserve the impact of important tokens as much as possible. Hence, we use the maximum dot product between the query and the key cache cluster. Specifically, given a query $\mathbf{q}\in\mathbb{R}^{1\times d}$, a key cache cluster of $s$ tokens $\mathbf{K}_{c}\in\mathbb{R}^{s\times d}$, the similarity can be designed as
$\mathrm{Sim}(\mathbf{q},\mathbf{K}_{c})=\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}\cdot\mathbf{k}=\max_{\mathbf{k}\in\mathbf{K}_{c}}\sum_{i=0}^{d-1}\mathbf{q}_{i}\mathbf{k}_{i}\leq\sum_{i=0}^{d-1}\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}_{i}\mathbf{k}_{i}.$ (1)
We obtain the upper bound of similarity, and further have
$\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}_{i}\mathbf{k}_{i}=\begin{cases}\mathbf{q}_{i}\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}&\mathrm{if}~\mathbf{q}_{i}\geq 0,\\ \mathbf{q}_{i}\min_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}&\mathrm{if}~\mathbf{q}_{i}<0.\end{cases}$
Define $\mathbf{r}^{\max}$ and $\mathbf{r}^{\min}$, where $\mathbf{r}_{i}^{\max}=\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}$ and $\mathbf{r}_{i}^{\min}=\min_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{k}_{i}$. Then, we have
$\mathrm{Sim}(\mathbf{q},\mathbf{K}_{c})\leq\sum_{i=0}^{d-1}\max(\mathbf{q}_{i}\mathbf{r}^{\max}_{i},\mathbf{q}_{i}\mathbf{r}^{\min}_{i}).$ (2)
Protocol complexity analysis. During the decoding stage, $\mathbf{r}_{i}^{\max}$ and $\mathbf{r}_{i}^{\min}$ of each cluster only need to be computed once. Hence, the computation cost can be amortized and become negligible. However, for each decoding step, we still need to compute $\mathcal{O}(LCd)$ multiplications, i.e., $\mathbf{q}_{i}\mathbf{r}^{\max}_{i}$ and $\mathbf{q}_{i}\mathbf{r}^{\min}_{i}$, as well as $\mathcal{O}(LCd)$ $\max$ operations in Equation (2), which still incur non-negligible overhead.
Table 2: The complexity analysis of token gathering protocol where $k_{1}=0.25T,k_{2}=0.25C$.
Bit Width # Comparison Lat. Comm. Example Lat. Example Comm.
Baseline Protocol $\log T$ $T$ $\mathcal{O}(T\log T)$ $\mathcal{O}(k_{1}T\log T)$ 4.780s 416.0MB
MPCache $\log C$ $C$ $\mathcal{O}(C\log C)$ $\mathcal{O}(k_{2}C\log C)$ 0.065s 1.125MB
[…]
Linearization and reordering. To avoid the MPC-unfriendly $\max$ operation in Equation (2), we further propose to approximate the similarity as below:
$\mathrm{Sim}(\mathbf{q},\mathbf{K}_{c})\approx\sum_{i=0}^{d-1}\alpha\cdot\mathbf{q}_{i}\mathbf{r}^{\max}_{i}+(1-\alpha)\cdot\mathbf{q}_{i}\mathbf{r}^{\min}_{i},$
where $\alpha\in[0,1]$ is a hyperparameter. As can be observed, when $\alpha=1$, $\mathbf{q}_{i}\mathbf{r}^{\max}_{i}$ is always selected while $\mathbf{q}_{i}\mathbf{r}^{\min}_{i}$ is always selected when $\alpha=0$. After the linearization, there is an opportunity to further reduce the multiplications by reordering the computation as
$\sum_{i=0}^{d-1}\alpha\cdot\mathbf{q}_{i}\mathbf{r}^{\max}_{i}+(1-\alpha)\cdot\mathbf{q}_{i}\mathbf{r}^{\min}_{i}=\sum_{i=0}^{d-1}\mathbf{q}_{i}\cdot(\alpha\mathbf{r}^{\max}_{i}+(1-\alpha)\mathbf{r}^{\min}_{i}).$ (3)
[…]
Figure 7: Comparison among maximum dot product (MP), average, and our method with different $\alpha$’s on TriviaQA.
Hierarchical KV cache clustering. Another question is “how to build the KV cache cluster efficiently?” Since larger cluster sizes have higher efficiency at the cost of worse performance, we propose to trade off the selection overhead and performance. Inspired by hierarchical reinforcement learning (xu2023haven, ), we propose to cluster the KV cache of adjacent tokens with a hierarchical structure as shown in Figure 6(a) that performs coarse-grained (larger cluster size) to fine-grained (smaller cluster size) selection progressively. Generally, we divide the KV cache into $n$ levels and progressively select the clusters level by level. At the fine-grained level, we only need to select from the remaining clusters, thereby reducing the selection complexity. Hierarchical structure, including the cluster size and selection ratios at each level, can influence the performance-efficiency trade-off. We discuss the trade-off in Section 5.4 and setups in Appendix F.1.
Figure 8: Commonality score between two adjacent layers.
4.4 Cross-layer Index-sharing Strategy
Based on the observation that adjacent layers share similar top-$k$ ranking of KV cache, we propose a cross-layer index-sharing strategy that enables adjacent layers to share the same selected token indices to further reduce the cost of dynamic selection. Since two adjacent layers show the highest commonality score in Figure 3(c), we choose to share the indices between two adjacent layers. In Figure 8, we observe the first two layers have a low commonality score while other layers have higher scores due to the residual, so we do not apply sharing to the first two layers. Cross-layer index-sharing effectively reduces the extra overhead introduced by dynamic selection. We discuss how the number of adjacent layers affects the trade-off in Section 5.4.
4.5 Security Analysis of MPCache
MPCache is built upon established cryptographic building blocks from 2PC lu2023bumblebee and 3PC dong2023puma that were already proven secure. Our introduced top-k protocol follows hou2023ciphergpt . Since our other optimizations do not introduce new protocols, but combine them to support efficient KV cache eviction, MPCache follows naturally from the security of these building blocks. MPCache assumes all parties are aware of the model architecture and number of pruned tokens, which is consistent with zhang2024heprune ; li2024seesaw ; kundu2023learning . We argue that it does not compromise the client’s data, nor does it enable the client to access the model’s parameters. Also, the shared indices between layers are ciphertexts, which won’t reveal any private information.
Figure 9: Comparison with prior-art fixed-pattern and static eviction algorithms on LongBench.
5 Empirical Evaluation
5.1 Experimental Setups
Models, datasets, and baselines. Our experiments are conducted on LLaMA-2, LongChat-7B-V1.5-32K, and LLaMA-3.1-8B-Instruct on LongBench (bai2023longbench, ), XSUM (narayan2018don, ), and Needle-in-a-Haystack kamradt2024llmtest . Refer to Appendix F.1 for details. We choose prior-art static and dynamic KV cache eviction baselines, including H2O (zhang2024h2o, ), StreamingLLM (xiao2023efficient, ), TOVA (oren2024transformers, ), SnapKV (li2024snapkv, ), InfLLM (xiao2024infllm, ), and LongCache (liu2024farewell, ). We leave the details of baselines to Appendix F.2.
Table 3: Comparison with prior-art dynamic KV cache eviction algorithms with different budgets.
Dataset
Cache Budget
InfLLM LongCache MPCache (ours)
[…]
HotpotQA Full 31.16 75.52 31.16 75.52 31.16 75.52
[…]
TriviaQA Full 82.67 75.52 82.67 75.52 82.67 75.52
[…]
NarrQA Full 17.02 75.52 17.02 75.52 17.02 75.52
[…]
PassageR Full 32.50 75.52 32.50 75.52 32.50 75.52
[…]
Qasper Full 27.58 75.52 27.58 75.52 27.58 75.52
[…]
Table 4: Evaluation on XSUM task.
Budget 10% 5%
Model $uparrow 7B\uparrow 13B\uparrow$
Full Cache 11.90 13.60 11.90 13.60
H2O 10.50 13.24 4.886 9.081
MPCache (ours) 11.10 13.44 10.08 13.08
Experimental setups. For performance evaluation, our experiments are conducted based on LongBench on an NVIDIA A100 80GB GPU. For efficiency evaluation, we use Secretflow (SPU V0.9.1) (ma2023secretflow, ) and follow the 3PC protocols of PUMA (dong2023puma, ). The latency is evaluated under the LAN setup. We evaluate the efficiency using the architecture of GPT-2 and LLaMA-2. We emphasize that the eviction ratio is determined based on the expected sequence length and the latency that can be tolerated by the server and the client in a realistic scenario. We explain the details of baselines, setups, and how to securely determine the hyper-parameters in Appendix F.1.
5.2 Performance Evaluation
[…]
Table 5: Extension to LLaMA-3.1-8B-Instruct on LongBench with an average KV cache budget of 2048 and 1024. The best result is highlighted in bold and the second best in underline.
Budget
Method
NarrativeQA
Qasper
HotpotQA
2WikiMQA
MuSiQue
TREC
SAMSum
TriviaQA
QMSum
PR-en
MF-en
2048 Full Cache 30.21 45.52 55.53 46.71 31.34 72.50 43.86 91.74 25.20 99.50 54.94
StreamingLLM 27.40 30.77 49.23 44.66 24.31 67.50 42.49 90.98 21.67 87.00 37.85
DuoAttention 25.61 42.31 52.36 42.14 28.17 66.00 43.36 89.93 22.11 98.50 55.53
SnapKV 28.53 39.13 54.32 46.59 28.48 55.55 43.10 92.04 23.92 99.50 53.91
Quest 28.00 45.55 53.73 44.76 29.82 68.56 44.05 90.90 24.91 99.50 55.40
MPCache 30.17 46.11 55.21 46.61 30.49 69.50 43.67 91.53 24.83 99.50 53.41
1024 Full Cache 30.21 45.52 55.53 46.71 31.34 72.50 43.86 91.74 25.20 99.50 54.94
StreamingLLM 26.64 27.48 47.31 42.03 24.17 63.50 42.76 88.84 21.31 88.00 35.59
DuoAttention 24.12 30.41 45.34 39.45 24.49 45.53 42.54 76.36 21.24 98.50 41.11
SnapKV 26.51 33.14 54.42 43.34 27.74 43.01 40.52 90.02 23.10 99.50 49.10
Quest 15.45 44.42 48.90 40.81 22.93 57.03 37.90 70.00 21.04 99.00 49.84
MPCache 29.47 46.20 55.38 46.55 31.14 69.50 43.02 91.84 24.77 99.50 52.70
Figure 10: Evaluation on per-token generation latency and communication compared with LongCache liu2024farewell , head merging rathee2024mpc , and ArkVale chenarkvale . Note that 5% and 10% mean the final KV cache budget.
Figure 11: Step-by-step ablation study.
Level 1 Coarse-grained Level 2 Fine-grained F1 Score (%) Comm. (MB) $s32(0.9)$ $s16(0.22)$ 29.6 163.5 $s32(0.7)$ $s16(0.28)$ 30.1 144.0 $s32(0.5)$ $s16(0.40)$ 30.2 140.2 $s32(0.3)$ $s16(0.67)$ 29.2 108.8 $s64(0.9)$ $s16(0.22)$ 29.5 158.1 $s64(0.7)$ $s16(0.28)$ 29.3 110.1 $s64(0.5)$ $s16(0.40)$ 29.1 104.9 $s64(0.3)$ $s16(0.67)$ 29.0 69.12 Table 6: Comparisons among different hierarchical structures.
5.3 Efficiency Evaluation
[…]
5.4 Ablation Studies of MPCache
[…]
Figure 12: Effect of the number of adjacent layers.
Effect of hierarchical structures. To trade off the model performance and dynamic selection overhead, we evaluate different hierarchical structures with a dynamic selection ratio of 20%. We choose different cluster sizes $s$ and selection ratios at different levels. From Table LABEL:tab:ablation_hierar, we find that when $s$ increases or the coarse-grained selection ratio decreases, the overhead decreases while the performance exhibits a downward trend. Moreover, appropriate coarse-grained selection may help improve the overall performance, e.g., the ratio changes from 90% to 50% with $s=32$;
Effect of the number of adjacent layers for cross-layer index-sharing. In response to Section 3, we evaluate the efficiency-performance trade-off of the number of adjacent layers on HotpotQA in Figure 12. As observed, when the number of adjacent layers increases, the latency is reduced while the performance degradation.
Benchmark on Needle-in-a-Haystack. We benchmark our similarity approximation algorithm on Needle-in-a-Haystack as shown in Figure 13. MPCache demonstrates comparable performance with full KV cache and outstanding performance compared with StreamingLLM.
Figure 13: Needle-in-a-Haystack. The x-axis denotes the length of the context (“haystack”) and the y-axis indicates the position where the “needle” (a short prompt) is inserted within the context.
Additional results. We present the influence of $\alpha$, discussion on 2PC protocol, comparison with average-based similarity in Appendix F.
6 Limitation and Conclusion
MPCache focuses on MPC inference, which still incurs non-negligible overhead compared to the plaintext counterpart. For certain tasks, such as those requiring the model to recall information from tokens considered unimportant, static eviction may fall short, which is a direction for future research.
In this work, we propose an MPC-friendly KV cache eviction framework named MPCache, that enables accurate and efficient private LLM inference. MPCache systematically combines static eviction and dynamic selection. To reduce the overhead of dynamic selection, we propose a series of MPC-friendly optimizations, including efficient similarity approximation, hierarchical KV cache clustering, and cross-layer index-sharing. Extensive evaluations demonstrate that MPCache consistently outperforms prior-art KV cache eviction baselines across different generation tasks and significantly reduces both latency and communication.
Acknowledgements
This work was supported in part by NSFC under Grant 62495102, Grant 92464104, and Grant 62341407, in part by Beijing Municipal Science and Technology Program under Grant Z241100004224015, and in part by Ant Group through CCF-Ant Research Fund.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Refer to the sections of Abstract and Introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Refer to the section of Limitation.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Refer to Appendix G.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Refer to the section of Experiments.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Refer to the code here.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe the detailed setups and the determination of hyperparameters in Appendix F.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We conduct experiments on multiple models and datasets, which require significant computational resources. In addition, we provide open access to the data and code, making it easy to reproduce the results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Our compute resources can be found in the Appendix F.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: /
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: /
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: /
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: /
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: /
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: /
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: /
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Detailed Background and Related Works
A.1 Private LLM Inference
Recently, private LLM inference has attracted an increasing amount of research attention. Iron [27] uses coefficient encoding to compute homomorphic encryption (HE)-based convolution layers efficiently and uses MPC to compute non-linear layers. PUMA [19] proposes a series of 3PC protocols for both linear and non-linear functions to support private LLM inference, even under the scale of LLaMA-7B. BumbleBee [53] proposes HE-based protocols that enable the multiplication of large matrices and efficient protocols for non-linear functions similar to PUMA. CipherGPT [31] uses subfield vector oblivious linear evaluation (sVOLE) to reduce the communication of MatMuls significantly. BOLT [60] proposes a baby-step giant-step (BSGS) strategy that reduces the number of rotations on ciphertexts. SIGMA [25] achieves private GPT inference with function secret sharing (FSS) and accelerates the computation on GPUs. PermLLM [94] proposes an efficient protocol for non-linear functions based on the random permutation. However, they mainly focus on protocol optimization, and still incur significant overhead, especially on long sequences.
There are also works directly replacing expensive non-linear functions in Transformers, e.g., Softmax and GeLU with MPC-friendly operations. For instance, MPCFormer [40] and Secformer [54] simplify Softmax by replacing exponential with an MPC-friendly quadratic function. MPCViT [84] replaces exponential in Softmax with ReLU and selectively uses scaling attention using neural architecture search (NAS). SAL-ViT [90] introduces external attention while PriViT [18] uses squaring function to replace Softmax. RNA-ViT [7] and Power-Softmax [96] use high-order polynomial approximation. [48] directly uniformly replaces GeLU with ReLU and replaces exponential in Softmax with ReLU. THE-X [9] approximates Softmax using an estimation model. Although these methods achieve high inference efficiency, they cannot avoid finetuning or re-training to preserve the model performance, making LLM development impractical. In conclusion, the above works still suffer from heavy overhead. Moreover, these works always handle full-length contexts during the LLM generation, incurring large latency and communication when handling long sequences. To solve this problem, our work aims to compress the KV cache in attention without training. Our proposed method can also be applied to different protocol frameworks for efficiency improvement.
A.2 KV Cache Compression
When tackling the LLM generation tasks, especially in long-context scenarios, the KV cache in the attention module becomes the most significant bottleneck due to the increasing sequence length. Therefore, how to effectively reduce the size of the KV cache is a high priority. System-level optimizations such as FlashAttention [14], FlashAttention-2 [13], FlashAttention-3 [65], and PagedAttention [39] have been proposed to alleviate the problem. Meanwhile, many recent research efforts have been devoted to algorithm-level optimizations. For example, quantization methods [30, 87, 35, 28, 51] have been proposed to compress KV cache to $1\sim 4$ bits, low-rank decomposition methods [64, 6, 92, 85] project the KV cache into low-rank space. In this work, we follow FlashAttention [14] to save the GPU memory and focus on another research line of algorithm-level optimization called KV cache eviction, which is designed to reduce the number of tokens and enable sparse attention without extra training.
KV cache eviction can be mainly categorized into 3 classes: 1) fixed-pattern algorithm: the position of important tokens is pre-defined before inference and remains consistent across decoding steps. However, this algorithm is not flexible for different LLMs and contexts [76, 3, 83]; 2) static algorithm: tokens are statically discarded and cannot be recovered in the subsequent decoding steps. This algorithm is usually efficient but suffers from significant performance degradation when the compression ratio is high [91, 50, 22, 42, 89, 78, 93, 10, 71, 33, 58]; 3) dynamic algorithm: tokens are dynamically selected across different decoding steps. This algorithm is much more flexible but the dynamic selection usually involves more expensive operations [74, 47, 67, 8]. We quantitatively compare existing methods in Table 1.
Here, we introduce recent works of KV cache eviction. StreamingLLM [76] proposes to keep a few initial tokens along with the recent tokens to recover the long-context performance. RazorAttention [66] theoretically analyzes the scope of effective attention vision for each head. Scissorhands [50], H2O [91], ALISA [93], spAtten [72], SnapKV [42] and LOOK-M [71], and TOVA [58] use the accumulated attention score of the historical tokens to select a small subset of KV cache. FastGen [22] and MInference 1.0 [33] propose to allocate different eviction policies for different heads based on the sparsity pattern of the prompt. PyramidKV [89], PyramidInfer [78], and SqueezeAttention [73] consider allocating different KV cache budgets for different layers. InfLLM [74] and LongCache [47] propose to dynamically select tokens based on the relationship between the current query and the key cache of previous tokens. RetrievalAttention [44] establishes connections from the query to its nearest keys and the decoding query can first search its nearest query and then obtain the most relevant key vectors. LazyLLM [20] introduces an aux cache to enable selective KV cache eviction. Keyformer [1] finds that the distribution after token pruning becomes uneven and proposes to smooth the distribution. Squeezed Attention [29] and ClusterKV [45] proposes to cluster KV cache based on the semantic information. HashAttention [16] and HashEvict [46] recently design hash functions to enable efficient KV cache eviction. To get rid of the dependency of similarity computation, SirLLM [81] uses token entropy while [17] uses the L2-norm of the key cache to measure the token importance.
However, these works are not designed or optimized for MPC since they either statically discard tokens that cause significant performance degradation or dynamically select tokens that introduce more complex and MPC-unfriendly operations. Recently, learning-based methods [75, 21] have also emerged, which require training to automatically explore the sparsity patterns and thus, are not the focus of this work.
A.3 Generative LLM Inference in Autoregressive-style
The generative LLM inference procedure is generally in an autoregressive style such as GPT-2 [61] and LLaMA [70], and mainly consists of two stages: 1) the prefill (prompt) stage and 2) the decoding (generation) stage.
Prefill stage. The prefill stage serves as the first step of generation. LLM takes a prompt sequence as input and generates key-value (KV) caches for each layer. The attention can be computed as
$\mathbf{O}_{\texttt{prompt}}=\mathrm{Softmax}(\mathbf{Q}_{\texttt{prompt}}\cdot\mathbf{K}_{\texttt{prompt}}^{\top}/\sqrt{d})\cdot\mathbf{V}_{\texttt{prompt}},$ (4)
where $\mathbf{Q}_{\texttt{prompt}}\in\mathbb{R}^{H\times T\times d}$ denotes the query and $\mathbf{K}_{\texttt{prompt}}\in\mathbb{R}^{H\times T\times d},\mathbf{V}_{\texttt{prompt}}\in\mathbb{R}^{H\times T\times d}$ denote the key and value cache, respectively. After the prefill stage, the KV cache is generated as $\mathbf{K}_{\texttt{cache}}\leftarrow\mathbf{K}_{\texttt{prompt}}$ and $\mathbf{V}_{\texttt{cache}}\leftarrow\mathbf{V}_{\texttt{prompt}}$. KV cache retains previously computed key-value pairs, eliminating the need for costly re-computation of previous key and value vectors [59]. Note that each layer is equipped with its unique KV cache and the generated KV cache is the foundation for the dowmstreaming decoding stage.
Decoding stage. The decoding stage uses and updates the stored KV cache to generate new tokens step-by-step. First, the KV cache is updated by concatenating new $\mathbf{k}\in\mathbb{R}^{H\times 1\times d}$ and $\mathbf{v}\in\mathbb{R}^{H\times 1\times d}$ as
$\mathbf{K}_{\texttt{cache}}\leftarrow\left[\mathbf{K}_{\texttt{cache}}||\mathbf{k}\right],~\mathbf{V}_{\texttt{cache}}\leftarrow\left[\mathbf{V}_{\texttt{cache}}||\mathbf{v}\right],$ (5)
where $\left[\cdot||\cdot\right]$ denotes tensor concatenation. Therefore, the attention can be computed as
$\mathbf{o}_{\texttt{dec}}=\mathrm{Softmax}(\mathbf{q}_{\texttt{dec}}\cdot\mathbf{K}_{\texttt{cache}}^{\top}/\sqrt{d})\cdot\mathbf{V}_{\texttt{cache}},$ (6)
where $\mathbf{q}_{\texttt{dec}}\in\mathbb{R}^{1\times d}$ denotes the current query. The attention output $\mathbf{o}_{\texttt{dec}}\in\mathbb{R}^{1\times d}$ is then sent to the multi-layer perceptron (MLP) layer for the subsequent computation.
Explanation about the use of KV cache. Although previous works overlook the use of KV cache [19, 53, 52], KV cache is already a fundamental component for existing generative LLMs to guarantee the inference correctness. If we do not use KV cache, the model has to re-compute all the previous tokens at each decoding step, significantly increasing the overhead. Hence, we apply KV cache to private LLM inference by default, and our aim is to reduce the computation associated with KV cache size for better efficiency. We also show the generation efficiency without KV cache in Section 5 and Appendix F.
Appendix B MPC Protocol Descriptions
B.1 Threat Model and Security
Consistent with previous works [56, 40, 19], MPCache adopts an honest-but-curious (a.k.a., semi-honest) security model in honest-majority [43] where parties follow the protocol specifications but may also try to learn more from the information than allowed. In our threat model, we assume all the parties are aware of the LLM architecture and number of pruned tokens, which is consistent with HEPrune [88], Seesaw [41], SENet [38], SNL [12], etc. We argue that this information does not compromise the client’s data or inference results, nor does it enable the client to access the model’s parameters.
B.2 2PC Protocol
We follow the 2PC protocols proposed in BumbleBee [53]. The protocols are built based on the 2-out-of-2 additive secret sharing (SS), where secret value $x\in 2^{\ell}$ is shared by two random values $x_{0},x_{1}\in 2^{\ell}$ such that $x=x_{0}+x_{1}\pmod{2^{\ell}}$, and party $P_{i}$ gets $x_{i}$ (denoted as [[x]][\![x]\!]). SS supports both addition and multiplication on the secret shares. Without special declaration, we compute in $2^{\ell}$ and omit $\pmod{2^{\ell}}$ for brevity. In the case of $\ell>1$ (e.g., $\ell=64$) which support arithmetic operations (e.g., ++, -, and $\cdot$), we refer to this type as arithmetic secret sharing. Boolean secret sharing refers to $\ell=1$ where (+,)(+,-) and $\cdot$ are replaced by bit-wise $\oplus$ and $\land$, respectively.
  • Addition. [[x+y]][\![x+y]\!] can be computed as $(x_{0}+y_{0},x_{1}+y_{1})$, where $P_{i}$ can compute its share locally.
  • Multiplication. We write the multiplication of two shared values as $xy]\!]=(x_{0}+x_{1})(y_{0}+y_{1})=x_{0}y_{0}+x_{1}y_{1}+x_{0}y_{1}+x_{1}y_{0}$ where two cross terms $x_{0}y_{1},x_{1}y_{0}$ can be computed using HE.
[53] uses HE scheme that is based on ring learning-with-error (RLWE). For more details about the 2PC protocol, please refer to [53, 55].
B.3 3PC Protocol
We follow the 3PC protocols proposed in PUMA [19]. The protocols are built based on the 2-out-of-3 replicated secret sharing (RSS), where a secret value $x\in 2^{\ell}$ is shared by three random values $x_{0},x_{1},x_{2}\in 2^{\ell}$ such that $x=x_{0}+x_{1}+x_{2}\pmod{2^{\ell}}$, and party $P_{i}$ gets $(x_{i},x_{i+1})$ (denoted as [[x]][\![x]\!]).
Let $(c_{1}$, $c_{2}$, $c_{3})$ be public constants, and ([[x]],[[y]])([\![x]\!],[\![y]\!]) be two secret-shared values. The secure addition and multiplication procedures are as follows:
  • Addition. $c_{1}x+c_{2}y+c_{3}]\!]$ can be computed as $(c_{1}x_{0}+c_{2}y_{0}+c_{3},c_{1}x_{1}+c_{2}y_{1},c_{1}x_{2}+c_{2}y_{2})$, where $P_{i}$ can compute its share locally. When $(c_{1}=1,c_{2}=1,c_{3}=0)$, we get [[x+y]][\![x+y]\!].
  • Multiplication. Parties follow steps: i) first, $P_{i}$ computes $z_{i}=x_{i}y_{i}+x_{i+1}y_{i}+x_{i}y_{i+1}$ locally; ii) parties then perform re-sharing by letting $P_{i}$ sends $z_{i}^{\prime}=\alpha_{i}+z_{i}$ to $P_{i-1}$, where $\alpha_{0}+\alpha_{1}+\alpha_{2}=0$ ($P_{i}$ can generate $\alpha_{i}$ using pseudorandom generators with negligible overhead as [56]); iii) finally, $\{(z_{0}^{\prime},z_{1}^{\prime}),(z_{1}^{\prime},z_{2}^{\prime}),(z_{2}^{\prime},z_{0}^{\prime})\}$ form the 2-out-of-3 replicated secret shares of [[xy]][\![xy]\!].
For more details about the 3PC protocol, please refer to [56, 55].
B.4 Token Gathering Protocol
Token gathering is used to retrieve tokens in the KV cache based on the indices, which has the same functionality as $\mathrm{torch.gather(tensor,indices)}$ in PyTorch programming. We illustrate the overall procedure in Figure 14.
For brevity, in Algorithm 2, we show the pipeline that retrieves one token from the key cache (we also omit the head dimension for simplification). The first step is converting a ciphertext index $\mathbf{id}]\!]$ into a ciphertext one-hot vector $\mathbf{o}]\!]\in\mathbb{R}^{1\times T}$ based on equal protocol $\Pi_{\mathrm{Equal}}$, where $T$ denotes the number of tokens. Given a ciphertext key cache $\mathbf{K}]\!]\in\mathbb{R}^{T\times d}$, where $d$ denotes the embedding dimension, multiplying $\mathbf{o}]\!]$ with $\mathbf{K}]\!]$ (matrix-vector multiplication $\Pi_{\mathrm{MatVec}}$) can generate an output with a dimension of $1\times d$, which is the retrieved token. To extend the case to retrieve $m$ tokens, we concatenate $m$ one-hot vectors to form a matrix $\mathbf{O}]\!]\in\mathbb{R}^{m\times T}$, and then multiply $\mathbf{O}]\!]$ with $\mathbf{K}]\!]$ (matrix-matrix multiplication $\Pi_{\mathrm{MatMul}}$) to generate an output with a dimension of $m\times d$, which is the retrieved $m$ tokens.
Note that token gathering protocol is also used on the value cache, and its indices are consistent with that of the key cache.
Figure 14: Illustration of token gathering procedure where we give an example of retrieving 3 tokens from 6 tokens.
1
Input : A ciphertext key cache $\mathbf{K}]\!]\in\mathbb{R}^{T\times d}$ and a ciphertext index $\mathbf{id}]\!]$.
Output : Key cache $\mathbf{K}]\!]^{\prime}\in\mathbb{R}^{1\times d}$ with the selected token.
2
31ex
4for $i\in[0,\dots,T-1]$ do
5   Parties jointly generate the one-hot vector as $\mathbf{o}[i]]\!]=\Pi_{\mathrm{Equal}}([\![\mathbf{id}]\!],i)$;
6
7Parties jointly compute the retrieved key cache as $\mathbf{K}]\!]^{\prime}=\Pi_{\mathrm{MatVec}}([\![\mathbf{o}]\!],[\![\mathbf{K}]\!])$;
return $\mathbf{K}]\!]^{\prime}$.
Algorithm 2 Token gathering protocol $\Pi_{\mathrm{Gather}}$ for retrieving one token
Appendix C Observation from Pattern Discovery of Large Attention Maps
It is sufficient to use a few tokens within the observation window to distinguish the attention patterns since the structure of attention maps is stable at different decoding steps [50, 78, 22, 42]. In Figure 15, we visualize the large attention map with hundreds of tokens on the PiQA [4] dataset to further verify our observation in Section 3. As can be observed, there are three types of tokens defined in Section 3: 1) IA tokens in red blocks which usually appear as attention sinks mentioned in StreamingLLM [76]. 2) IC tokens in orange blocks; 3) UIA tokens in blue blocks; The pattern of attention maps motivates us to statically discard the UIA tokens which may have negligible impact on further generation, and dynamically select important tokens from IC tokens at each decoding step for sparse attention computation.
Figure 15: Attention patterns across different layers on LLaMA-7B.
Appendix D Pseudocode of MPCache Framework
We describe the overall flow of our MPCache in detail as shown in Algorithm 3.
1
Input :  Input sequence prompt; LLM model $\mathcal{M}$; number of layers and attention heads $L$ and $H$; dynamic selection ration $\eta\in[0,1]$; three types of tokens IA, IC, and UIA (introduced in Section 3); cluster size $s$; decoding steps $E$.
Output : Attention output and evicted KV cache.
2
31ex
4Step 1: Look-once static eviction during prefill stage
5for $l\in[0,\dots,L-1]$ do
6  $\mathbf{Q}^{(l)},\mathbf{K}^{(l)},\mathbf{V}^{(l)}\leftarrow\mathcal{M}(\texttt{prompt})$;
7 if $l\%2==0$ then
8     $\mathbf{ATTN}_{\texttt{window}}^{(l)}\leftarrow\mathrm{Softmax}(\mathbf{Q}^{(l)}[:,-len(\texttt{prompt})\times 0.2:,:]\cdot\mathbf{K}^{(l)\top})$;
9    $\{IA,IC,UIA\}^{(l)}\leftarrow\mathrm{static\_evict}(\mathbf{ATTN}_{\texttt{window}}^{(l)})$;
10 else
11     $\{IA,IC,UIA\}^{(l)}\leftarrow\{IA,IC,UIA\}^{(l-1)}$; $\triangleright$ Cross-layer index sharing (Section 4.4)
12
13 $\mathbf{K}^{(l)}\leftarrow\mathrm{token\_gather}(\mathbf{K}^{(l)},index=\{IA,IC\}^{(l)})$;
14 $\mathbf{V}^{(l)}\leftarrow\mathrm{token\_gather}(\mathbf{V}^{(l)},index=\{IA,IC\}^{(l)})$;
15
16Step 2: Query-aware dynamic selection during decoding stage
17for $e\in[0,\dots,E-1]$ do
18  for $l\in[0,\dots,L-1]$ do
19
20    if $l\%2==0$ then
21
22       $\mathbf{sim}\leftarrow\mathrm{Sim}(\mathbf{q}^{(l,e)},\mathbf{K}^{(l,e)},cluster\_{size}=s)$; $\triangleright$ Follow Equation (3)
23       $\mathbf{index}^{(l,e)}\leftarrow\mathrm{topk}(\mathbf{sim},k=\mathbf{K}^{(l,e)}.size()\times\eta)$;
24    else
25        $\mathbf{index}^{(l,e)}\leftarrow\mathbf{index}^{(l-1,e)}$; $\triangleright$ Cross-layer index sharing (Section 4.4)
26
27    $\mathbf{K}^{(l,e)}\leftarrow\mathrm{token\_gather}(\mathbf{K}^{(l,e)},index=\mathbf{index}^{(l,e)})$;
28    $\mathbf{V}^{(l,e)}\leftarrow\mathrm{token\_gather}(\mathbf{V}^{(l,e)},index=\mathbf{index}^{(l,e)})$;
29    $\mathbf{O}^{(l,e)}\leftarrow\mathrm{Softmax}(\mathbf{q}^{(l,e)}\cdot\mathbf{K}^{(l,e)\top}/\sqrt{d})\cdot\mathbf{V}^{(l,e)}$; $\triangleright$ Sparse attention
30
31
return $\mathbf{O,K,V}$.
Algorithm 3 MPCache: KV cache eviction combining static and dynamic algorithm
Appendix E Paradigm Comparison with Dynamic Algorithm
We compare the paradigm between the dynamic algorithm and our MPCache as shown in Figure 16. Through performing static eviction during the prefill stage, we improve the efficiency of all decoding steps since the token selector (i.e., similarity approximation and top-$k$ ranking) only needs to handle a smaller number of tokens during the decoding stage (30% in this case).
Figure 16: Paradigm comparison between (a) dynamic algorithm and (b) our proposed MPCache combining static eviction and dynamic selection. MPCache discards unimportant tokens to reduce the decoding overhead (red texts mean the differences).
Appendix F Supplemental Experiments
F.1 Supplemental Setups
Experimental environment. The model performance is evaluated with LongBench on an NVIDIA A100 80GB GPU in PyTorch. The latency is evaluated with Secretflow under the LAN setup [63] with 377MBps bandwidth and 0.3ms echo latency [63] on Intel(R) Xeon(R) Gold 5220R CPU @ 2.20GHz. To save GPU memory and avoid the out-of-memory (OOM) error when processing long contexts, we leverage FlashAttention [14] during the prefill stage. Since securely evaluating a full-size 7B model in SPU exceeds our hardware resources, we set a smaller embedding dimension of 1024 in our evaluation.
Datasets. In our experiments, we use XSUM [57], LongBench [2], and Needle-in-a-Haystack [23]. LongBench is a benchmark for bilingual, multitask, and comprehensive assessment of long context understanding capabilities of large language models. We also use the XSUM [57] dataset which is used for the evaluation of abstractive single-document summarization systems. Needle-in-a-Haystack is an evaluation method that randomly inserts key information into long texts to form prompts for LLMs. The test aims to detect whether models can extract such key information from extensive texts, thereby assessing the models’ capabilities in processing and understanding long documents.
KV cache clustering configuration. For hierarchy, we in practice choose a two-level hierarchical structure, i.e., $n=2$, and when the final dynamic selection ratio $\alpha<0.5$, we drop 50% clusters at the 1st hierarchical level. For the XSUM dataset, we use a cluster size of 8 at the 1st hierarchical level and 4 at the 2nd hierarchical level. For the long-context LongBench, we use larger clusters, i.e., 32 at the 1st hierarchical level and 16 at the 2nd hierarchical level.
Static eviction. During the static eviction, we compute the accumulated attention sum only using the last 20% tokens in the prompt. However, using 20% tokens still incurs a CUDA out-of-memory (OOM) error when processing a long-context prompt (i.e., longer than 24k tokens). To solve this problem, we adaptively adjust the ratio to 10% tokens instead. We also notice that the choice of the ratio won’t cause significant performance fluctuations, so we omit the discussion in this paper.
How to securely determine the hyper-parameters, e.g., eviction ratios in the realistic scenario? In practice, we believe the static and dynamic eviction ratio should be determined based on two aspects: 1) the expected sequence length and 2) the latency that can be tolerated/afforded by the server and the client. Our framework proposes a novel algorithm with configurable hyper-parameters to enable exploring the Pareto front of the LLM performance and efficiency. The framework has been validated across different benchmarks and different LLMs, which demonstrates its value and potential for practical use cases.
F.2 Detailed Descriptions about Baselines
StreamingLLM [76] follows a fixed eviction pattern, i.e., keep the local tokens and initial tokens) across different decoding steps. H2O [91] and TOVA [58] statically prune the KV cache and these discarded tokens cannot be recovered at subsequent decoding steps. SnapKV [42] statically prunes the KV cache only during the prefill stage. InfLLM [74] employs block-level dynamic token selection during the decoding stage. It requires selecting several representative tokens within a block and computing the relevance score using these representative tokens. LongCache [47] uses the idea of cosine similarity between the query and key cache of previous tokens to select relevant tokens without static eviction and token clustering. Note that LongCache separates the positional embedding (PE) from the KV cache. Since our focus is on the dynamic selection metric in this work, we do not apply PE separation in LongCache.
F.3 Supplemental Ablation Studies
Effect of hyper-parameter $\alpha$. To study how $\alpha$ impacts the similarity approximation, we select different $\alpha$’s on different datasets as shown in Figure 18. As can be observed, although the effects of different $\alpha$ do not occur in a certain pattern, we can still discover some patterns related to the dataset from the trend in Figure 18: on TriviaQA, the model may prefer larger $\alpha$ while it may prefer smaller $\alpha$ on HotpotQA instead. Since $\alpha=0.6$ shows relatively better performance in these cases, we choose $\alpha=0.6$ by default in our experiments.
Figure 17: The influence trend of similarity approximation with different $\alpha$ values ranging from 0 to 1.
Figure 18: Comparison with average-based similarity approximation. MP means maximum dot product.
Figure 19: Extension MPCache to 2PC protocol.
Comparison with average-based similarity approximation. A straightforward and efficient way to aggregate the information of a key cache cluster is the average. We compare our proposed method ($\alpha=0.6$) with average-based similarity on the XSUM dataset with a cluster size of 16 in Figure 18. Specifically, we perform dynamic selection with different ratios after 75% tokens are statically discarded. As can be observed, using average suffers from significant performance degradation under different ratios. With the compression ratio increasing, the degradation of the average-based method becomes more serious. An intuitive explanation is that using the average of a cluster may make some important tokens averaged and ignored. In contrast, our approximation can effectively maintain the model performance. We theoretically analyze the similarity approximation algorithm in Appendix G.
[…]
Static eviction ratio. We show the ablation study to explain our choice of static eviction ratio in Table 7. We choose the static eviction ratio that does not significantly affect the performance on the datasets, and different static eviction ratios can influence the overall decoding performance.
Table 7: Model performance with different static eviction ratios on HotpotQA.
Eviction Ratio (%) F1 Score (%)
0 31.16
55 30.82
70 30.54
75 28.89
Appendix G Theoretical Analysis of Similarity Approximation
As mentioned in Section 2, given query $\mathbf{q}\in\mathbb{R}^{H\times 1\times d}$, key cache $\mathbf{K}\in\mathbb{R}^{H\times T\times d}$, and value cache $\mathbf{V}\in\mathbb{R}^{H\times T\times d}$, the overall goal of KV cache eviction is to find an optimal policy $\mathcal{P}$ to minimize the gap between the attention outputs (here we omit $\mathbf{V}$ for simplification) as
$\mathcal{P}^{*}=\mathrm{argmin}|\mathrm{Softmax}(\mathbf{q}\cdot\mathbf{K}^{\top})-\mathrm{Softmax}(\mathbf{q}\cdot\mathbf{K}^{\prime\top})|,$ (7)
where $\mathbf{K}^{\prime}$ is a subset of $\mathbf{K}$ selected by $\mathcal{P}$. However, when grouping $\mathbf{K}$ into clusters for efficiency, the problem becomes challenging. We denote the key cache cluster as $\mathbf{K}_{c}$ (cluster size is $s$), and our goal is to find a way to accurately approximate the similarity between $\mathbf{q}$ and the key cluster $\mathbf{K}_{c}$. This problem is equivalent to “how can we effectively aggregate the cluster information to obtain a cluster representation and measure its importance?”
We assume there exists a function $\phi$ that aggregates the key cluster $\mathbf{K}_{c}$, and we define the optimization problem as
$\min|\sum_{j=0}^{s-1}\exp(\mathbf{q}\cdot\mathbf{K}_{cj}^{\top})-\mathbf{q}\cdot\phi(\mathbf{K}_{c}^{\top})|.$ (8)
Drawback of average-based clustering. As mentioned, the simplest way to represent the cluster is the average and $\phi(\mathbf{K}_{cj}^{\top})$ becomes $\sum_{j=0}^{s-1}\mathbf{K}_{cj}^{\top}/s$. This happens to be equivalent to directly drop $\exp$ in $\sum_{j=0}^{s-1}\exp(\mathbf{q}\cdot\mathbf{K}_{cj}^{\top})$, introducing information loss. Intuitively, if there are tokens with low importance and tokens with high importance within one cluster, the overall importance will be averaged, leading to the neglect of the crucial tokens.
Different from the average-based method, using the max dot product aims to protect the crucial tokens with large scores as much as possible. This observation is aligned with [67].
$\mathrm{MaxDotProduct}:~~\mathbf{q}\cdot\phi(\mathbf{K}_{c}^{\top})=\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}\cdot\mathbf{k}.$ (9)
In order to approximate $\max_{\mathbf{k}\in\mathbf{K}_{c}}\mathbf{q}\cdot\mathbf{k}$ without accessing all the tokens in $\mathbf{K}_{c}$, we utilize the bounding volume proposed by [36]. Our optimizations are described in Section 4.3.
Appendix H Overall Secure Inference Framework
Take 2PC as an example in Figure 20, we illustrate the secure inference framework following [53] where the server owns the proprietary LLM parameter and the client possesses private input data. During inference, the data is secretly shared between two parties. Linear layers are computed using the HE protocol, and non-linear layers require interactive protocols between the two parties based on oblivious transfer (OT) and HE. Figure 20 illustrates the detailed data flow of MPCache during the token generation.
Figure 20: Overall secure inference data flow of MPCache during token generation.